### v1

# Where Did It All Go Wrong? A Hierarchical Look into Multi-Agent Error Attribution

Abstract
Error attribution in Large Language Model (LLM) multi-agent systems presents a significant challenge in debugging and improving collaborative AI systems. Current approaches to pinpointing agent and step level failures in interaction traces - whether using all-at-once evaluation, step-by-step analysis, or binary search - fall short when analyzing complex patterns, struggling with both accuracy and consistency. We present ECHO (Error attribution through Contextual Hierarchy and Objective consensus analysis), a novel algorithm that combines hierarchical context representation, objective analysis-based evaluation, and consensus voting to improve error attribution accuracy. Our approach leverages a positional-based leveling of contextual understanding while maintaining objective evaluation criteria, ultimately reaching conclusions through a consensus mechanism. Experimental results demonstrate that ECHO outperforms existing methods across various multi-agent interaction scenarios, showing particular strength in cases involving subtle reasoning errors and complex interdependencies. Our findings suggest that leveraging these concepts of structured, hierarchical context representation combined with consensus-based objective decision-making, provides a more robust framework for error attribution in multi-agent systems.
1 Introduction
The evolution of large language models (LLMs) has driven their implementation as collaborative, specialized agents working together in structured systems [1, 2]. These systems break down complex challenges into manageable components, distributing them among purpose-specific agents that work in concert toward common objectives [3, 4]. Such coordination demands effective orchestration, with each agent performing specialized functions. Building these collaborative systems requires careful consideration of interconnected graph structures—mapping agent relationships, information flows, and logical constraints. These structural elements form the foundation for expandable and flexible multi-agent frameworks [5].
Multi-Agent Systems (MASs) have demonstrated remarkable performance across coding [6], medical QA [7], and financial decision-making [8]. However, their multi-step nature makes them vulnerable to compounding errors, where early mistakes amplify through subsequent steps and can derail the entire system. As a result, identifying the initial error’s source—both agent and step—becomes crucial for mitigating such failures and improving these systems.
As MASs grow in complexity, manual error attribution becomes unscalable, necessitating an automated approach. However, according to the Who&When benchmark [9], even SOTA LLMs — both closed-source (GPT-4o [10], o1 [11]) and open-source (Llama 4 [12]) — struggle with this task. The complexity stems from interdependent agent interactions, large context sizes, and the need to understand both local and global context within interaction traces. Traditional debugging approaches falter in these dynamic systems where errors are often subtle and context-dependent.
Automated error attribution in LLM-based MAS has explored varying approaches to failure log analysis. For example, all-at-once methods expose LLMs to complete logs simultaneously for agent and step identification [9]. Alternatively, step-by-step approaches evaluate interactions sequentially until detecting an error [9]. More sophisticated binary search methods iteratively narrow the search space by having LLMs determine which half of the trace contains the critical mistake [9].
This paper presents ECHO (Error attribution through Contextual Hierarchy and Objective consensus analysis), a novel approach to error attribution in multi-agent systems, that addresses these limitations, by guiding error attribution through developing a hierarchical context representation of the entire interaction trace, providing independent objective analyses across these contexts and cross-validating their findings via consensus voting.
2 Related Work
2.1 Evaluation of LLM Agents
Much research has focused on automated LLM evaluation across various tasks [13, 14, 15], with recent efforts shifting toward assessing the agentic capabilities of LLMs. AgentBench [16] tests task completion across diverse environments like online shopping and householding, while MLAgentBench [17] evaluates LLM agents’ abilities to perform machine learning experimentation. AssistantBench [18] features realistic, time-intensive web tasks requiring multi-turn interactions. Agent-as-a-Judge [19] leverages LLM agents to evaluate their other agent peers through tool interactions. ToolFormer [20], AnyTool [21], and ToolBench [22] benchmark LLMs’ API-calling capabilities via tool-use. As LLM agent interactions evolve, evaluation frameworks now address multi-agent scenarios: MultiAgentBench [23] measures coordination and competitive dynamics, while SwarmBench [24] assesses swarm intelligence through coordination tasks like foraging and flocking.
2.2 Error Attribution
While LLM evaluation typically measures task completion and output quality, error attribution focuses on systematically classifying error types and patterns to understand fundamental failure modes [25, 26, 27, 28]. ReaLMistake [29] assesses LLMs’ error attribution capabilities in generated output, while SynCheck [30] leverages decoding dynamics to verify sentence trustworthiness and backtracks the unfaithfulness of generated sentences. Self-Backtracking [31] trains LLMs to localize and correct their own reasoning errors through supervised fine-tuning. Similarly, Process Reward Models (PRMs) [32] evaluate intermediate reasoning step correctness through process annotation, while ProcessBench [33] measures PRMs’ step-wise evaluation capabilities. These approaches focus solely on single-agent error analysis, whereas the Who&When dataset [9] extends error attribution to multi-agent settings by identifying both the responsible agent and erroneous step.
3 ECHO Methodology
Error attribution in multi-agent systems demands three fundamental capabilities: context understanding to capture interaction patterns, error analysis to detect failure points, and decision synthesis to determine final attribution. ECHO addresses these through hierarchical context representation, decoupled objective analysis (at both agent and step levels), and confidence-weighted consensus voting, each targeting specific attribution challenges.
Figure 1: Figure 1: ECHO Architecture. The system comprises: (1) Hierarchical Context - processes traces through 4 compression layers (L1-L4: full content → milestones) with specialized modules for handoffs, decisions, errors, and patterns; (2) Decoupled Analysis - uses 6 specialized agents (conservative to balanced) generating structured outputs with evidence, confidence scores, and hypotheses; (3) Consensus Voting - aggregates analyses via confidence-weighted voting and disagreement resolution.
3.1 Hierarchical Context Representation
Error attribution in multi-agent systems faces a fundamental tension: while comprehensive context is crucial for accurate attribution, processing limitations make it impractical to analyze full interaction traces $\tau$ of $n$ agents with equal detail. Traditional approaches typically restrict analysis to immediate positional neighbors (±1 agent), presenting critical limitations. This narrow window fails to capture long-range dependencies where errors propagate across multiple steps and misses crucial context from causally relevant agents outside this window (e.g. error in initial problem formulation manifesting several steps later, or early incorrect assumptions propagating through multiple agents before causing a failure). ECHO addresses these limitations through a multi-layered hierarchical context representation ($C$) that captures both local and global interaction patterns.
Hierarchical representation operates at 4 levels $L_{1}$ through $L_{4}$, as seen in Appendix A.1 and A.3, to extract key information from agent / step interactions (via regex pattern matching). While LLMs could potentially be used for more flexible content extraction, the computational overhead would be significant - a trade-off we explore further in Section 4.3. Its implementation employs specialized content extraction mechanisms for each layer $C_{i}$, as seen below:
  1. 1.
    Immediate Context layer ($L_{1}$): Encompasses the target agent $i$ and its direct neighbors ($\tau_{i\pm 1}$), preserving complete reasoning chains and interaction patterns. This layer maintains full agent content, including detailed decision-making processes, input-output relationships, and intermediate computations. This granular preservation enables precise analysis of local decision points and immediate error manifestations. Examples of these are: "Let me analyze the data… [Terminal Output]: DataFrame loaded, showing 500 rows".
  2. 2.
    Local Context Layer ($L_{2}$): Spans $\tau_{i\pm 2,3}$ steps from the target agent, focusing on tactical decision sequences and their interconnections. By preserving key reasoning chains while filtering routine operations, this layer reveals short-range error propagation patterns and local decision dependencies. This intermediate scope bridges the gap between immediate interactions and broader strategic patterns. Examples of these are focusing on specific patterns such as conclusive statements (e.g., "I conclude the optimal parameters are x=0.5, y=1.0") and logical transitions (e.g., "Therefore, we should use gradient descent").
  3. 3.
    Distant Context Layer ($L_{3}$): Covers $\tau_{i\pm 4,5,6}$ steps through strategic compression, distilling agent interactions into concise outcome summaries. This layer captures essential state transitions, critical assumptions, and warning signals while filtering peripheral details. This compression enables identification of longer-range dependencies without overwhelming the analysis with excessive detail. Examples of these are critical state changes (e.g., "Model training failed: insufficient data"), error conditions (e.g., "Classification blocked: missing features"), and handoff information (e.g., "Received dataset with normalized features").
  4. 4.
    Global Context Layer ($L_{4}$): Encompasses the remaining interaction trace ($\tau_{remainder}$) through highly compressed milestone-based representation. By retaining only strategically significant decision points and major state transitions, this layer enables system-wide consistency checking and identification of broad error patterns. This high-level view ensures local decisions align with global objectives while maintaining computational feasibility. Examples of these include error propagation signals (e.g., "Persistent data quality errors"), major state transitions (e.g., "Switched from classification to regression"), and cross-agent dependencies (e.g., "Reusing parameters from Agent 2’s optimization").
This graduated extraction and compression approach adapts to different context types (handoff, decision quality, error propagation, general), enabling context-aware information preservation at each layer. Lastly, it is worth noting that when using regex-based extraction, real-world systems can be designed to generate reasoning traces with extractable keywords, effectively pre-seeding content for regex. This maintains efficiency while achieving accuracy comparable to more complex methods, making it practical for real applications.
3.2 Objective Analysis
ECHO employs a panel of $k$ objective analysis agents that evaluate the full interaction trace $\tau$ through hierarchical context $C$ (algorithm in Appendix A.1). Each analysis agent independently assesses all steps and provides confidence scores $\sigma_{j}$, enabling nuanced attributions that can identify distributed responsibility, interaction effects, and systemic issues spanning multiple agents or steps.
The objective analysis framework leverages the hierarchical context through conversation summaries that preserve contextual relationships, enabling multi-scale error attribution. Agents examine errors across all layers: detailed decisions ($L_{1}$), state transitions ($L_{2}$), error propagation ($L_{3}$), and systemic issues ($L_{4}$). Furthermore, attribution is restricted to steps with explicit reasoning, ensuring precise analysis across context scales while maintaining step-level granularity.
The effectiveness of objective analysis depends critically on agent diversity to mitigate systematic biases. Rather than deploying identical agents that might propagate the same analytical biases, ECHO employs a panel of specialized analysts with roles $\rho_{j}$ (as seen in the code in Appendix A.5):
  1. 1.
    Conservative Analyst: Requires strong, clear evidence for attribution, prefers single-agent attributions, maintains high confidence thresholds, and only attributes with definitive proof
  2. 2.
    Liberal Analyst: Considers multi-agent error scenarios, identifies subtle error patterns, accepts moderate confidence thresholds, and makes attributions with reasonable evidence
  3. 3.
    Detail-Focused Analyst: Examines specific evidence and exact wording, identifies subtle inconsistencies, focuses on fine-grained analysis, and prioritizes concrete evidence over patterns
  4. 4.
    Pattern-Focused Analyst: Recognizes broader reasoning chains, tracks error propagation patterns, identifies recurring themes, and analyzes overall reasoning structure
  5. 5.
    Skeptical Analyst: Questions underlying assumptions, explores alternative explanations, challenges conventional attributions, and examines validity of ground truth
  6. 6.
    General Analyst: Maintains balanced perspective, considers all evidence types equally, focuses on obvious, impactful errors, and provides baseline objective evaluation
Each analysis agent follows a structured evaluation protocol, examining both individual steps and their hierarchical context relationships. The agents’ analyses ($\epsilon_{j}$) are guided by specialized prompts that reflect their assigned analytical focus, producing outputs with four components: (1) investigation summary, (2) detailed evaluations with error likelihood scores [0,1] and evidence, (3) primary conclusion specifying attribution type, responsible agent(s), mistake step, and confidence score $\sigma_{j}$, and (4) alternative hypotheses. This structure ensures consistent, comparable outputs for consensus voting across diverse analytical perspectives.
This approach prevents echo-chamber effects in attribution through intentional perspective diversity. While this was implemented in ECHO through role-specific prompts and temperatures, the framework accommodates other diversification methods (different models, LLM configurations, or reasoning strategies) - all serving the core principle of bias mitigation through analytical diversity.
3.3 Consensus Voting
The final component of ECHO is a consensus voting mechanism that aggregates analyses from the panel of $k$ objective analysts, as seen in Appendix A.7. The voting system operates through weighted confidence consensus, where each analyst’s attribution is weighted by their reported confidence level $\sigma_{j}$, subject to a minimum confidence threshold $\delta$ to filter out low-confidence attributions.
The consensus mechanism processes 3 key components from each analysis $A_{t}^{j}$ (where $t\in\{\text{agent},\text{step}\}$): (1) Primary Conclusions - that considers core attribution decisions including the attribution type, responsible agent(s), and specific error step; (2) Agent Evaluations - that provides assessments of each agent’s error likelihood and supporting evidence; and (3) Alternative Hypotheses - that considers secondary attribution possibilities that maintain reasonable likelihood.
The voting process follows a hierarchical decision structure, that first determines the winning conclusion type through weighted confidence aggregation of each analyzer’s agent and step vote ($V_{a},V_{s}$). Then for single/multi-agent conclusions, ECHO identifies specific agents through confidence-weighted attribution votes ($\omega_{a}$), after which it validates and aggregates step-level predictions ($\omega_{s}$). Finally it synthesizes supporting reasoning from the highest-confidence attributions.
The system explicitly handles disagreements through disagreement analysis $\phi$ by examining: (1) conclusion diversity (number of different conclusion types); (2) confidence spread across analysts ($\max(\sigma_{j})-\min(\sigma_{j})$); (3) attribution consistency between agent and step-level predictions; (4) need for additional review when high disagreement exists (spread > 0.5) or multiple conflicting high-confidence attributions emerge.
This structured voting approach ensures that the final attribution (ConsensusResult($\omega_{a},\omega_{s},\phi$)) reflects the collective analysis while maintaining sensitivity to strong minority opinions and potential edge cases. The combination of confidence weighting and explicit disagreement handling provides robustness against individual analyst biases while preserving valuable insights from dissenting views.
4 Results and Analysis
4.1 Experimental Setup
We evaluate our implementations using the Who&When benchmark [9], a comprehensive dataset comprising failure logs from LLM-powered multi-agent systems. This benchmark is particularly suitable for error attribution evaluation as it provides fine-grained failure annotations across diverse scenarios, including both algorithm-generated and hand-crafted multi-agent interactions.
Dataset Characteristics - The benchmark consists of 2 distinct subsets: (1) Algorithm-Generated: systematic failure logs from automated multi-agent interactions; (2) Hand-Crafted: curated failure scenarios designed to capture complex error patterns. Each failure log is annotated with the failure-responsible agent; the specific step where the error occurred; and detailed failure reasoning.
Evaluation Protocol - We assess each implementation across 4 experimental conditions: (1) Algorithm-Generated with ground truth; (2) Algorithm-Generated without ground truth; (3) Hand-Crafted with ground truth; and (4) Hand-Crafted without ground truth. Attribution accuracy is measured at 2 levels: agent-level (referring to the correct identification of the failure-responsible agent; and step-level (referring to the correct identification of the specific error step).
Implementation Details - Our framework employs Claude 3.5 Sonnet v2 as the base LLM, with varying temperatures between 0.3 and 0.9 across 3 randomly sampled analysis agents (from a pool of 6 specialists) for every scenario, to balance analytical diversity while reducing computational overhead. We use a confidence threshold ($\delta=0.3$) for filtering attributions in the consensus voting mechanism, to balance inclusion of diverse perspectives while filtering out highly uncertain predictions. This implementation assumes position-based contextual importance, independence between analysis agents, and treats errors as primarily additive rather than emergent.
4.2 Comparative Analysis of Implementations
We evaluate 4 progressive implementations of error attribution, each building upon the limitations of its predecessor:
Implementation 1 (I1) - Fixed Context Window: Represents the baseline approach using a fixed context window (±1 step), where a context-aware agent analyzes each step with its immediate neighbors, followed by a final judge agent for attribution decisions.
Implementation 2 (I2) - Hierarchical Context: Enhances I1 by replacing the fixed context window with a four-layer hierarchical context representation, maintaining the same attribution mechanism but providing graduated access to the full interaction trace.
Implementation 3 (I3) - Objective Analysis: Builds upon I2’s hierarchical context by replacing the context-aware agent and judge agent with a panel of specialized objective analyst agents and performing consensus voting on those outcomes, enabling diverse analytical perspectives.
Implementation 4 (I4) - Decoupled Attribution: Refines I3 by separating the attribution process into two distinct phases: agent-level attribution to identify responsible agents, followed by step-level attribution to pinpoint specific error points, allowing for more focused analysis at each level.
4.2.1 Performance of ECHO
Table 1: Performance of ECHO across different datasets and configurations
Hand-Crafted Dataset Algorithm-Generated Dataset
Method With GT Without GT With GT Without GT P-value
Agent-Level Accuracy
Random 0.120 0.120 0.291 0.291 <0.001
All-at-Once 0.577 0.529 0.563 0.530 0.032
Step-by-Step 0.360 0.343 0.397 0.283 <0.001
Binary Search 0.517 0.362 0.441 0.301 0.007
ECHO (ours) 0.684 0.679 0.688 0.672 -
Step-Level Accuracy (Exact)
Random 0.042 0.042 0.191 0.191 <0.001
All-at-Once 0.060 0.021 0.152 0.145 <0.001
Step-by-Step 0.066 0.069 0.274 0.178 0.003
Binary Search 0.069 0.069 0.240 0.166 0.012
ECHO (ours) 0.281 0.268 0.288 0.272 -
Step-Level with Tolerance (Hand-Crafted with GT)
All-at-Once Step-by-Step Binary Search ECHO (ours) P-value
±1 step 0.149 0.166 0.138 0.351 <0.001
±2 steps 0.223 0.189 0.190 0.386 0.004
±3 steps 0.350 0.209 0.224 0.421 0.029
±4 steps 0.380 0.294 0.319 0.579 0.006
±5 steps 0.428 0.351 0.362 0.614 0.008
Token Cost (Hand-Crafted with GT)
All-at-Once Step-by-Step Binary Search ECHO (ours) -
Tokens 17,106 87,720 34,659 53,701 -
P-values compare ECHO against each baseline using chi-squared test.
ECHO demonstrates robust statistically significant performance (P < 0.05) across both hand-crafted and algorithm-generated datasets, with several notable patterns emerging from the results:
Agent-Level Attribution:
The system achieves consistent agent-level accuracy ($\sim$68%) across all configurations, with minimal degradation (1–2%) when ground truth is withheld. Furthermore, the similar performance across both dataset types (hand-crafted: 68.4%, algorithm-generated: 68.8%) further demonstrates ECHO’s robust generalization capabilities across different interaction patterns. This ground truth independence, combined with consistent cross-dataset performance, makes ECHO particularly valuable for deployment in dynamic, real-world AI systems where there is generally an absence of labeled data.
Step-Level Attribution: Although ECHO still outperforms prior methods with statistical significance, step-level precision still proves challenging, with exact matches at 27–28%. However, accuracy improves significantly when accounting for step tolerance. With ±3 steps, accuracy reaches 42.1% for hand-crafted dataset with ground truth; and when further extended to ±5 steps, accuracy reaches 61.4% - again, maintaining superior performance over baselines.
Impact of Ground Truth: The minimal difference in performance between configurations with and without ground truth (typically <2% variation) suggests that ECHO’s architecture effectively leverages interaction patterns and context for attribution, rather than relying on ground truth information. This is particularly important for real-world applications where ground truth may not be available.
Token Counts: ECHO demonstrates moderate token efficiency, using $\sim$54K tokens compared to $\sim$88K for Step-by-Step and $\sim$17K for All-at-Once approaches on hand-crafted datasets. While higher than the simplest baseline, ECHO’s token usage reflects its balanced approach to comprehensive analysis, maintaining reasonable processing costs of \$0.15 per example with Claude 3.5 Sonnet v2.
The results demonstrate ECHO’s capability to provide reliable agent-level attribution while offering flexible step-level identification with tolerance ranges. The system’s consistent performance across different datasets and configurations, combined with its reasonable computational overhead, makes it suitable for practical applications in error attribution for multi-agent systems.
4.3 Impact of ECHO Components Via Ablation
Table 2: Ablation Study: Impact of Each Component
Hand-Crafted Dataset Algorithm-Generated Dataset
Implementation With GT Without GT With GT Without GT P-value
Agent-Level Accuracy
Fixed Context (I1) 0.286 0.265 0.461 0.452 -
+ Hierarchical (I2) 0.447 0.429 0.523 0.508 0.037
+ Objective Analysis (I3) 0.610 0.589 0.651 0.635 0.043
+ Decoupled Attribution (I4) 0.684 0.679 0.688 0.672 0.196
Step-Level Accuracy
Fixed Context (I1) 0.151 0.143 0.157 0.140 -
+ Hierarchical (I2) 0.170 0.166 0.192 0.175 0.398
+ Objective Analysis (I3) 0.232 0.218 0.461 0.444 <0.001
+ Decoupled Attribution (I4) 0.281 0.268 0.288 0.272 0.211
Token Cost
Fixed Context (I1) 4.02M 3.93M 319K 317K -
+ Hierarchical (I2) 7.70M 7.66M 407K 405K -
+ Objective Analysis (I3) 67.5K 66.5K 12.6K 12.5K -
+ Decoupled Attribution (I4) 33.0K 32.5K 12.8K 12.7K -
Hand-Crafted Dataset results for I1 and I2 based on limited sample of shorter traces
P-values compare each component with previous implementation
The Question of Unifying or Decoupling Objective Analyses
The analysis of context length reveals key insights about unified versus decoupled attribution (Table 2). For shorter algorithm-generated traces, unified analysis achieves strong results (65.1% agent-level, 46.1% step-level accuracy), while decoupling shows mixed effects: marginal improvement in agent-level accuracy (68.8%) but notable decline in step-level precision (28.8%). This suggests that splitting attribution tasks can be counterproductive when operating within the LLM’s comfortable processing range ($\sim$13K tokens unified vs. $\sim$6K+7K tokens decoupled).
For longer hand-crafted traces, unified analysis struggles with step-level precision (23.2%) while maintaining moderate agent-level accuracy (61.0%). Decoupling improves performance (68.4% agent-level, 28.1% step-level) while significantly reducing per-task tokens (from $\sim$67K to $\sim$13K+19K). This reveals context-length dependency: shorter traces benefit from unified analysis’s complete context, while longer traces near model limits favor decoupled analysis’ complexity management.
The Computational Overhead Resolution By Using Objective Analysis
The shift to objective analysis (I3) reveals key limitations of context-aware agents in error attribution. While fixed (I1) and hierarchical (I2) context implementations leverage trace information, their repeated analysis proves computationally intensive. Objective analysis dramatically improves efficiency, reducing token usage by 60-110x for hand-crafted cases, and by 25-30x for algorithm-generated cases, while still improving accuracy ($\sim$+16.3% agent-level and $\sim$+6.2% step-level at P < 0.05) as shown in Table 2. These results demonstrate that objective analysis is crucial for practical deployment, enabling efficient processing of longer traces while maintaining comprehensive context benefits.
Switching from Fixed-Window Context to Hierarchical Context
Hierarchical context shows clear advantages over fixed context windows within the context-aware framework. Shifting from fixed ±1 step (I1) to hierarchical context (I2) yields significant improvements in hand-crafted datasets: +16.1% in agent-level accuracy (P < 0.05) and +1.9% in step-level accuracy (not significant), as shown in Table 2. A similar trend is exhibited for the algorithm-generated dataset, with a +6.2% in agent-level accuracy and +3.8% in step-level accuracy. This validates graduated context preservation, where detail decreases with distance from target steps. While analysis for the hand-crafted dataset was limited to shorter traces due to computational constraints with context-aware agents, the consistent accuracy improvements highlight hierarchical context’s value - a benefit fully realized when combined with objective analysis.
Table 3: Additional Ablations on Hand-Crafted Dataset with Ground Truth
Implementation Agent-Level Step-Level Token Cost P-value
LLM Capability
Claude-3 Haiku 0.618 0.164 60,008 0.042
ECHO (Claude-3.5 Sonnet) 0.684 0.281 53,701 -
Claude-3.7 Sonnet 0.788 0.269 57,223 0.036
Hierarchical Context Method
ECHO (Regex-based) 0.684 0.281 53,701 -
LLM-based 0.750 0.438 165,932 -
Objective Analyst Panel Size
ECHO (3 Analysts) 0.684 0.281 53,701 -
6 Analysts 0.667 0.259 87,603 0.483
Tested on limited sample of shorter traces due to computational constraints
P-values compare with ECHO baseline for each sections
Additional Ablation Studies
Additional ablation studies on the hand-crafted dataset with ground truth reveal key design insights. Using stronger reasoning models significantly improve performance (P < 0.05): Claude-3.7 Sonnet achieves 78.8% agent-level and 26.9% step-level accuracy versus Haiku’s 61.8% and 16.4%, maintaining similar token usage ($\sim$60K). Switching from regex-based context extraction to LLM-based context extraction shows promise (75.0% agent-level, 43.8% step-level) but proves computationally intractable ($\sim$166K vs. 54K tokens, with extraction alone using $\sim$145K). Lastly, expanding the analyst panel size from 3 to 6 yields minimal gains (68.4% vs. 66.7% agent-level) while doubling token usage (88K vs. 54K), validating our default configuration’s efficiency. These results have been shown in Table 3.
5 Conclusion
We introduce ECHO, a novel approach to error attribution in multi-agent systems that combines hierarchical context representation, decoupled objective analysis, and confidence-weighted consensus voting. Our results demonstrate substantial improvements over existing baseline methods, particularly for longer traces where traditional approaches become prohibitive. Beyond immediate applications, ECHO’s precise error attribution capabilities have broader implications for AI development: in reinforcement learning, it can help identify and eliminate false steps that may reduce model efficacy, while in single-agent optimization, it enables targeted prompt refinement for systems like GEPA [34] and DEEVO [35]. Future directions include developing relevance-based architectures for dynamic context preservation, enhancing the consensus mechanism through multi-agent debate protocols [35], and incorporating error severity metrics as well as partial correctness evaluation. Additionally, the Who&When benchmark could be expanded with standardized problem-type categorization and complex architectural patterns to better reflect real-world scenarios. As multi-agent systems proliferate, ECHO’s efficient context handling and bias mitigation approach provides a crucial foundation for both debugging current systems and advancing AI development practices.
Appendix A Appendix
A.1 ECHO Algorithm
Algorithm 1 ECHO: Error Attribution through Contextual Hierarchy and Objective Consensus Analysis
0:
[…]
4:$k$ : number of analysis agents
[…]
5: Attribution of error to specific agent(s) and step(s)
6:Procedure HierarchicalContextExtraction($\tau$):
[…]
8:for each agent $i\in\{1,...,n\}$ do
9:  $L_{1}\leftarrow$ ExtractFullContext($\tau_{i\pm 1}$)
10:  $L_{2}\leftarrow$ ExtractKeyDecisions($\tau_{i\pm 2,3}$)
11:  $L_{3}\leftarrow$ CompressSummaries($\tau_{i\pm 4,5,6}$)
12:  $L_{4}\leftarrow$ ExtractMilestones($\tau_{remainder}$)
13:  $C_{i}\leftarrow\{L_{1},L_{2},L_{3},L_{4}\}$
14:end for
15:return $C$
16:Procedure DecoupledAgentAndStepAnalysis($C,\alpha,\gamma$):
17:for type $t\in\{\text{agent},\text{step}\}$ do
18:  $A_{t}\leftarrow\emptyset$ {Init results}
19:  for each analyst $j\in\{1,...,k\}$ do
20:   $\rho_{j}\leftarrow$ AnalystRole($j$)
21:   if $\gamma\neq\text{None}$ then
22:    $\epsilon_{j}\leftarrow$ Eval($t,C,\rho_{j},\gamma$)
23:   else
24:    $\epsilon_{j}\leftarrow$ Eval($t,C,\rho_{j}$)
25:   end if
26:   $\sigma_{j}\leftarrow$ ConfidenceScore($\epsilon_{j}$)
27:   $A_{t}^{j}\leftarrow\{\epsilon_{j},\sigma_{j}\}$
28:  end for
29:end for
30:return $(A_{\text{agent}},A_{\text{step}})$
31:Procedure ConsensusVoting($A_{a},A_{s},\delta$):
[…]
33:for each analysis pair $(A_{a}^{j},A_{s}^{j})$ do
34:  if $\sigma_{j}\geq\delta$ then
35:   $V_{a},V_{s}\leftarrow V_{a}\cup\{A_{a}^{j}\},V_{s}\cup\{A_{s}^{j}\}$
36:  end if
37:end for
[…]
40:return ConsensusResult($\omega_{a},\omega_{s},\phi$)
[…]
43:return ConsensusVoting($A_{a},A_{s},\delta$)
A.2 Fixed-Window Context
def extract_agent_contexts(
conversation_history: List[Dict[str, Any]]
) -> List[Tuple[Optional[Dict[str, Any]], Dict[str, Any], Optional[Dict[str, Any]]]]:
␣␣␣␣Extractagentcontextsfromconversationhistory.
␣␣␣␣Eachcontextincludesthepreviousagent,currentagent,andnextagent.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣conversation_history:Listofconversationturnswithagentinformation
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Listoftuplescontaining(prev_agent,current_agent,next_agent)foreachagent
contexts = []
for i in range(len(conversation_history)):
# Get previous agent (None if first agent)
prev_agent = conversation_history[i - 1] if i > 0 else None
# Get current agent
current_agent = conversation_history[i]
# Get next agent (None if last agent)
next_agent = conversation_history[i + 1] if i < len(conversation_history) - 1 else None
contexts.append((prev_agent, current_agent, next_agent))
return contexts
A.3 Hierarchical Context Extraction
def extract_key_decision(
agent_content: str, max_words: int = 50, context_type: str = "decision_quality"
) -> str:
␣␣␣␣Extractkeydecisionormainpointfromagentcontentusingregexpatterns.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
␣␣␣␣␣␣␣␣max_words:Maximumwordsintheextractedkeydecision
␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Keydecisionormainpointfromtheagent’scontent
if not agent_content.strip():
return "Nocontentavailable"
if context_type == "handoff":
patterns = [
[…]
elif context_type == "decision_quality":
patterns = [
[…]
elif context_type == "error_propagation":
patterns = [
[…]
else: # general
patterns = [
r"(?:I(?:will|should|needto|decidedto|concludethat|believe|think|determine))([^.!?]*[.!?])",
r"(?:Therefore|Thus|So|Hence),?([^.!?]*[.!?])",
r"(?:Theanswer|Theresult|Thesolution)(?:is|appearstobe|seemstobe)([^.!?]*[.!?])",
r"Letme([^.!?]*[.!?])",
r"(?:Myapproach|Mystrategy|Myplan)(?:is|willbe)([^.!?]*[.!?])",
# Try to find pattern matches
for pattern in patterns:
matches = re.findall(pattern, agent_content, re.IGNORECASE)
if matches:
decision = matches[0].strip()
words = decision.split()[:max_words]
return "".join(words) + ("..." if len(decision.split()) > max_words else "")
# Fallback: take the first sentence or first max_words
sentences = agent_content.split(".")
if sentences:
first_sentence = sentences[0].strip()
if not first_sentence.endswith("."):
first_sentence += "."
words = first_sentence.split()[:max_words]
return "".join(words) + ("..." if len(first_sentence.split()) > max_words else "")
# Final fallback: just truncate
words = agent_content.split()[:max_words]
return "".join(words) + ("..." if len(agent_content.split()) > max_words else "")
def summarize_agent(agent_content: str, max_words: int = 20, context_type: str = "general") -> str:
␣␣␣␣Createabriefsummaryofagentcontentusingregexpatterns.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
␣␣␣␣␣␣␣␣max_words:Maximumwordsinthesummary
␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Briefsummaryoftheagent’scontent
if not agent_content.strip():
return "Nocontentavailable"
# Remove excessive whitespace and newlines
cleaned_content = "".join(agent_content.split())
if context_type == "handoff":
patterns = [
[…]
elif context_type == "decision_quality":
patterns = [
[…]
elif context_type == "error_propagation":
patterns = [
[…]
else: # general
patterns = [
r"(?:Inconclusion|Toconclude|Therefore|Thus|So|Hence),?([^.!?]*[.!?])",
r"(?:The(?:answer|result|solution|output))(?:is|appearstobe|seemstobe)([^.!?]*[.!?])",
r"(?:I(?:found|determined|concluded|calculated))([^.!?]*[.!?])",
# Try pattern matching first
for pattern in patterns:
matches = re.findall(pattern, cleaned_content, re.IGNORECASE)
if matches:
summary = matches[0].strip()
words = summary.split()[:max_words]
return "".join(words) + ("..." if len(summary.split()) > max_words else "")
# Fallback: take first sentence and truncate
sentences = cleaned_content.split(".")
if sentences:
first_sentence = sentences[0].strip()
words = first_sentence.split()[:max_words]
return "".join(words) + ("..." if len(first_sentence.split()) > max_words else "")
# Final fallback
words = cleaned_content.split()[:max_words]
return "".join(words) + ("..." if len(cleaned_content.split()) > max_words else "")
def obtain_milestones(agent_content: str, max_words: int = 15, context_type: str = "general") -> str:
␣␣␣␣Extractmilestone-basedinformationfromagentcontentusingregexpatterns.
␣␣␣␣Thisprovidesahigherlevelofabstractionthanbriefsummariesfordistantcontexts.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
␣␣␣␣␣␣␣␣max_words:Maximumwordsintheextractedmilestones
␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Milestone-basedinformationfromtheagent’scontent
if not agent_content.strip():
return "Nomilestonesavailable"
# Remove excessive whitespace and newlines
cleaned_content = "".join(agent_content.split())
if context_type == "handoff":
patterns = [
[…]
elif context_type == "decision_quality":
patterns = [
[…]
elif context_type == "error_propagation":
patterns = [
[…]
else: # general
patterns = [
[…]
# Try to find pattern matches
for pattern in patterns:
matches = re.findall(pattern, cleaned_content, re.IGNORECASE)
if matches:
milestone = matches[0].strip()
words = milestone.split()[:max_words]
return "".join(words) + ("..." if len(milestone.split()) > max_words else "")
# Fallback: extract first meaningful sentence
sentences = cleaned_content.split(".")
if sentences:
first_sentence = sentences[0].strip()
words = first_sentence.split()[:max_words]
return "".join(words) + ("..." if len(first_sentence.split()) > max_words else "")
# Final fallback
words = cleaned_content.split()[:max_words]
return "".join(words) + ("..." if len(cleaned_content.split()) > max_words else "")
def extract_agent_contexts_hierarchical(
conversation_history: List[Dict[str, Any]], dataset_name: str = ""
) -> List[Dict[str, Any]]:
␣␣␣␣Extracthierarchicalagentcontextsfromconversationhistory.
␣␣␣␣Usesgraduateddetaillevelsbasedondistancefromcurrentagent.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣conversation_history:Listofconversationturnswithagentinformation
␣␣␣␣␣␣␣␣dataset_name:Nameofthedatasetbeingprocessed(affectshowagentinfoisextracted)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Listofdictionariescontaininghierarchicalcontextforeachagent
contexts = []
for current_idx in range(len(merged_history)):
current_agent = conversation_history[current_idx]
# Build hierarchical context for this agent
hierarchical_context = {
"current_agent": current_agent,
"context_levels": {
"immediate": [], # Distance 1: Full detail
"nearby": [], # Distance 2-3: Key decisions
"distant": [], # Distance 4-6: Brief summaries
"milestones": [], # Distance >6: Milestones
# Process all other agents based on their distance
for i, agent in enumerate(conversation_history):
if i == current_idx:
continue # Skip current agent
distance = abs(i - current_idx)
agent_info = {
"index": i,
"name": agent["name"],
"role": agent["role"],
"distance": distance,
if distance == 1: # Immediate context: Full detail
agent_info["content"] = agent["content"]
agent_info["detail_level"] = "full"
hierarchical_context["context_levels"]["immediate"].append(agent_info)
elif distance <= 3: # Nearby context: Key decisions
agent_info["content"] = extract_key_decision(agent["content"])
agent_info["detail_level"] = "key_decisions"
hierarchical_context["context_levels"]["nearby"].append(agent_info)
elif distance <= 6: # Distant context: Brief summaries
agent_info["content"] = summarize_agent(agent["content"])
agent_info["detail_level"] = "summary"
hierarchical_context["context_levels"]["distant"].append(agent_info)
else: # Milestone context: High-level milestones for very distant agents
agent_info["content"] = obtain_milestones(agent["content"])
agent_info["detail_level"] = "milestones"
hierarchical_context["context_levels"]["milestones"].append(agent_info)
# Sort each level by original conversation order
for level in hierarchical_context["context_levels"].values():
level.sort(key=lambda x: x["index"])
contexts.append(hierarchical_context)
return contexts
A.4 Context Step Aware Agent
class ContextAwareStepAgent:
␣␣␣␣Context-AwareStepAgentthatanalyzesanagentinthecontextofitspreviousandnextagents
␣␣␣␣toarguewhytheerrorhappenedinthisagent’sstep.
def __init__(
self,
model_id: str = "us.anthropic.claude-3-5-sonnet-20241022-v2:0",
temperature: float = 1.0,
␣␣␣␣␣␣␣␣InitializetheContext-AwareStepAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration
self.bedrock_model = BedrockModel(
model_id=model_id,
temperature=temperature,
top_p=0.9,
max_tokens=4096,
self.system_prompt = """
␣␣␣␣␣␣␣␣YouareaContext-AwareStepAgentthatanalyzesanagent’sactionsinthecontextofthepreviousandnextagents.
␣␣␣␣␣␣␣␣YourtaskistoarguewhytheerrorhappenedinYOURagent’sstep.
␣␣␣␣␣␣␣␣Yourtask:
␣␣␣␣␣␣␣␣1.Analyzewhatinformationwasreceivedbyyouragentfromthepreviousagent(ifany)
␣␣␣␣␣␣␣␣2.Analyzewhatinformationwasgeneratedbyyouragent
␣␣␣␣␣␣␣␣3.Analyzehowyouragent’soutputaffectedthenextagent(ifany)
␣␣␣␣␣␣␣␣4.MakeastrongargumentforwhyYOURAGENTcausedthefinalwronganswer,usingthegroundtruthasevidence
␣␣␣␣␣␣␣␣Input:
␣␣␣␣␣␣␣␣-GroundTruth:[GROUND_TRUTH]
␣␣␣␣␣␣␣␣-FinalAnswer:[FINAL_ANSWER]
␣␣␣␣␣␣␣␣-AgentContext:Informationabouttheprevious,current,andnextagents
␣␣␣␣␣␣␣␣Outputyourresponsewiththefollowingclearsectionheaders:
␣␣␣␣␣␣␣␣##Purpose:
␣␣␣␣␣␣␣␣Describethepurposeofthisagentstep-whatwasthisagenttryingtoaccomplish?
␣␣␣␣␣␣␣␣##AssumptionsandInformation:
␣␣␣␣␣␣␣␣Listtheassumptionsandinformationthisagentwasgivenfromthepreviousagentorcontext.
␣␣␣␣␣␣␣␣##Errors:
␣␣␣␣␣␣␣␣Describewhatthisagentdidwrong(ifanything).Bespecificaboutanymistakes,misunderstandings,orincorrectreasoning.
␣␣␣␣␣␣␣␣##Evidence:
␣␣␣␣␣␣␣␣Provideevidencefromthegroundtruththatsupportsyourerrorattribution.Explainhowthisagent’sactionsdirectlyledtothewrongfinalanswer.
␣␣␣␣␣␣␣␣Remember:YoumustarguethatYOURagentcausedtheerror.Bepersuasiveanduseevidence.
self.agent = Agent(
system_prompt=self.system_prompt,
model=self.bedrock_model,
def analyze_agent(
self,
step_id: str,
prev_agent: Optional[Dict[str, Any]],
current_agent: Dict[str, Any],
next_agent: Optional[Dict[str, Any]],
ground_truth: Optional[str],
final_answer: str,
query: str = "",
) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Analyzeanagentinthecontextofitspreviousandnextagentsandgenerateanargument
␣␣␣␣␣␣␣␣forwhythisagentcausedtheerror.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣step_id:TheIDofthestep(e.g.,"step_1")
␣␣␣␣␣␣␣␣␣␣␣␣prev_agent:Thepreviousagent(orNoneiffirstagent)
␣␣␣␣␣␣␣␣␣␣␣␣current_agent:Thecurrentagentbeinganalyzed
␣␣␣␣␣␣␣␣␣␣␣␣next_agent:Thenextagent(orNoneiflastagent)
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣JSONargumentforwhythisagentcausedtheerror
prompt = f"""
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣{ground_truth_section}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣AgentContext:
␣␣␣␣␣␣␣␣{agent_context}
␣␣␣␣␣␣␣␣Pleaseanalyzethisagentinthecontextofthepreviousandnextagents,andprovideastrongargumentforwhyTHISagentcausedthefinalwronganswer.
␣␣␣␣␣␣␣␣Usethesectionheadersspecifiedinyourinstructions(Purpose,AssumptionsandInformation,Errors,Evidence).
agent_result = self.agent(prompt)
# Extract text from AgentResult
response_text = ""
if hasattr(agent_result, "message") and "content" in agent_result.message:
content = agent_result.message["content"]
if isinstance(content, list) and len(content) > 0 and "text" in content[0]:
response_text = content[0]["text"]
elif isinstance(content, str):
response_text = content
# Create a dictionary with the step_id, agent_name, and the full text response
result = {
"step_id": step_id,
"agent_name": current_agent["name"],
"analysis": response_text,
"token_usage": token_usage,
return result
def analyze_agent_hierarchical(
self,
step_id: str,
hierarchical_context: Dict[str, Any],
ground_truth: Optional[str],
final_answer: str,
query: str = "",
) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Analyzeanagentusinghierarchicalcontextandgenerateanargument
␣␣␣␣␣␣␣␣forwhythisagentcausedtheerror.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣step_id:TheIDofthestep(e.g.,"step_1")
␣␣␣␣␣␣␣␣␣␣␣␣hierarchical_context:Dictionarycontaininghierarchicalcontextinformation
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣JSONargumentforwhythisagentcausedtheerror
current_agent = hierarchical_context["current_agent"]
prompt = f"""
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣{ground_truth_section}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣AgentContext:
␣␣␣␣␣␣␣␣{agent_context}
␣␣␣␣␣␣␣␣Pleaseanalyzethisagentinthehierarchicalcontextoftheentireconversation,andprovideastrongargumentforwhyTHISagentcausedthefinalwronganswer.
␣␣␣␣␣␣␣␣Usethesectionheadersspecifiedinyourinstructions(Purpose,AssumptionsandInformation,Errors,Evidence).
␣␣␣␣␣␣␣␣Note:Younowhaveaccesstothefullconversationcontextatdifferentdetaillevels:
␣␣␣␣␣␣␣␣-Immediatecontext:Fulldetailsofadjacentagents
␣␣␣␣␣␣␣␣-Nearbycontext:Keydecisionsfromagents2-3stepsaway
␣␣␣␣␣␣␣␣-Distantcontext:Briefsummariesofagents4+stepsaway
␣␣␣␣␣␣␣␣Considerhowinformationanderrorsmighthavepropagatedacrosstheentireconversationwhenmakingyourargument.
agent_result = self.agent(prompt)
# Extract text from AgentResult
response_text = ""
if hasattr(agent_result, "message") and "content" in agent_result.message:
content = agent_result.message["content"]
if isinstance(content, list) and len(content) > 0 and "text" in content[0]:
response_text = content[0]["text"]
elif isinstance(content, str):
response_text = content
# Create a dictionary with the step_id, agent_name, and the full text response
result = {
"step_id": step_id,
"agent_name": current_agent["name"],
"analysis": response_text,
"context_type": "hierarchical",
"token_usage": token_usage,
return result
A.5 Objective Analysis Agent
class ObjectiveAnalysisAgent:
␣␣␣␣ObjectiveAnalysisAgentthatanalyzesallagentsinaconversationobjectively
␣␣␣␣todetermineerrorattributionwithoutforcedbias.
def __init__(
self,
model_id: str = "us.anthropic.claude-3-5-sonnet-20241022-v2:0",
temperature: float = 0.7,
analyst_focus: str = "general",
␣␣␣␣␣␣␣␣InitializetheObjectiveAnalysisAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration
self.bedrock_model = BedrockModel(
model_id=model_id,
temperature=temperature,
top_p=0.9,
max_tokens=4096,
# Create specialized system prompt based on analyst focus
focus_instructions = self._get_focus_instructions(analyst_focus)
self.system_prompt = f"""
␣␣␣␣␣␣␣␣YouareanObjectiveAnalysisAgentconductinganimpartialinvestigationtodetermineerrorattributioninamulti-agentconversation.
␣␣␣␣␣␣␣␣ANALYSTSPECIALIZATION:{focus_instructions}
␣␣␣␣␣␣␣␣Yourtask:
␣␣␣␣␣␣␣␣1.AnalyzeALLagentsintheconversationobjectively(notjustonespecificagent)
␣␣␣␣␣␣␣␣2.Determinewhichagent(s)mostlikelycausedthefinalwronganswer
␣␣␣␣␣␣␣␣3.Determinewhichstep/turnintheconversationthemistakeoccurred
␣␣␣␣␣␣␣␣4.Provideconfidencescoresandreasoningforyourconclusions
␣␣␣␣␣␣␣␣Youhaveaccesstohierarchicalcontextshowing:
␣␣␣␣␣␣␣␣-Immediateagents:Fulldetails
␣␣␣␣␣␣␣␣-Nearbyagents:Keydecisions
␣␣␣␣␣␣␣␣-Distantagents:Briefsummaries
␣␣␣␣␣␣␣␣Theagentsarenumberedsequentially(Agent1,Agent2,etc.)correspondingtotheirstep/turnindexintheconversation.
␣␣␣␣␣␣␣␣Possibleconclusions:
␣␣␣␣␣␣␣␣-Singleagenterror:Onespecificagentcausedthemistakeataspecificstep
␣␣␣␣␣␣␣␣-Multi-agenterror:Multipleagentscontributedtothemistakeacrossspecificsteps
␣␣␣␣␣␣␣␣OutputyourresponseasvalidJSONwrappedin<json></json>tags:
␣␣␣␣␣␣␣␣<json>
␣␣␣␣␣␣␣␣␣␣"analysis_summary":"Brief overview of your investigation approach and findings",
␣␣␣␣␣␣␣␣␣␣"agent_evaluations":[
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"agent_name":"agent_name",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"step_index":1,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"error_likelihood":0.0-1.0,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Why this agent may or may not have caused the error",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"evidence":"Specific evidence supporting your assessment"
␣␣␣␣␣␣␣␣␣␣"primary_conclusion":{{
␣␣␣␣␣␣␣␣␣␣␣␣"type":"single_agent"|"multi_agent",
␣␣␣␣␣␣␣␣␣␣␣␣"attribution":["agent_name(s)"]ornull,
␣␣␣␣␣␣␣␣␣␣␣␣"mistake_step":1,
␣␣␣␣␣␣␣␣␣␣␣␣"confidence":0.0-1.0,
␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Detailed explanation of your primary conclusion including which step the error occurred"
␣␣␣␣␣␣␣␣␣␣"alternative_hypotheses":[
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"type":"conclusion_type",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"attribution":["agent_name(s)"]ornull,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"mistake_step":1,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"confidence":0.0-1.0,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Alternative explanation"
␣␣␣␣␣␣␣␣</json>
␣␣␣␣␣␣␣␣Bethorough,objective,andconsiderallpossibilitiesincludingthatnosingleagentmaybeclearlyatfault.
␣␣␣␣␣␣␣␣Payspecialattentiontoidentifyingthespecificstep/turnwheretheerroroccurred.
self.agent = Agent(
system_prompt=self.system_prompt,
model=self.bedrock_model,
def analyze_conversation(
self,
conversation_contexts: List[Dict[str, Any]],
ground_truth: Optional[str],
final_answer: str,
query: str = "",
conversation_history: Optional[List[Dict[str, Any]]] = None,
) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Analyzetheentireconversationobjectivelytodetermineerrorattribution.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣conversation_contexts:Listofhierarchicalcontextdictionariesforallagents
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Dictionarycontainingobjectiveanalysisresults
# Create a comprehensive context summary for analysis
context_summary = self._create_conversation_summary(conversation_history)
prompt = f"""
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣{ground_truth_section}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣ConversationAnalysis:
␣␣␣␣␣␣␣␣{context_summary}
␣␣␣␣␣␣␣␣Pleaseconductanobjectiveanalysisofthisconversationtodetermineerrorattribution.
␣␣␣␣␣␣␣␣Focusonidentifyingwhichspecificagent(s)causedtheerrorthatledtotheincorrectfinalanswer.
␣␣␣␣␣␣␣␣OutputyouranalysisintheJSONformatspecifiedinyourinstructions.
agent_result = self.agent(prompt)
# Extract text from AgentResult
response_text = ""
if hasattr(agent_result, "message") and "content" in agent_result.message:
content = agent_result.message["content"]
if isinstance(content, list) and len(content) > 0 and "text" in content[0]:
response_text = content[0]["text"]
elif isinstance(content, str):
response_text = content
try:
# Parse the JSON response
analysis_result = validate_json(response_text)
analysis_result["raw_response"] = response_text
# Add token usage to the result
if token_usage:
analysis_result["token_usage"] = token_usage
return analysis_result
except ValueError as e:
print(f"Errorparsingobjectiveanalysisresponse:{e}")
print(f"Rawresponse:{response_text}")
# Return a basic structure if parsing fails
return {
"analysis_summary": "Errorparsingresponse",
"agent_evaluations": [],
"primary_conclusion": {
"type": "single_agent",
"attribution": None,
"confidence": 0.0,
"reasoning": "Failedtoparseanalysisresponse",
"alternative_hypotheses": [],
"raw_response": response_text,
"token_usage": token_usage,
def _create_conversation_summary(
self, conversation_contexts: List[Dict[str, Any]]) -> str:
␣␣␣␣␣␣␣␣Createacomprehensivesummaryoftheconversationforobjectiveanalysis.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣conversation_contexts:Listofhierarchicalcontextdictionaries
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Formattedconversationsummary
summary = []
# Extract agent information from contexts with their ORIGINAL step indices
agents_info = []
step_indices = list(range(len(conversation_contexts)))
for i, context in enumerate(conversation_contexts):
current_agent = context["current_agent"]
agents_info.append(
"step_index": step_indices[i]
"name": current_agent["name"],
"role": current_agent["role"],
"content": current_agent["content"],
# Create structured summary with clear step indexing
summary.append("===CONVERSATIONAGENTS===")
for agent in agents_info:
summary.append(f"Step{agent[’step_index’]}-{agent[’name’]}({agent[’role’]}):")
summary.append(f"{agent[’content’]}")
summary.append("")
# Add context relationships for the first few agents as examples
summary.append("===HIERARCHICALCONTEXTEXAMPLE===")
if conversation_contexts:
sample_context = format_hierarchical_context(conversation_contexts[0])
summary.append("ContextstructureforAgent1(showinghierarchicaldetaillevels):")
summary.append(
sample_context[:1000] + "..." if len(sample_context) > 1000 else sample_context
[…]
def _get_focus_instructions(self, analyst_focus: str) -> str:
␣␣␣␣␣␣␣␣Getspecializedinstructionsbasedonanalystfocus.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣analyst_focus:Thetypeofanalystfocus
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Specializedinstructionsstring
focus_map = {
"conservative": "Youareaconservativeanalystwithhighconfidencethresholds.Onlyattributeerrorswhenyouhavestrong,clearevidence.Prefersingle-agentattributionsovermulti-agentones.Becautiousaboutmakingattributionswithoutdefinitiveproof.",
"liberal": "Youarealiberalanalystmorewillingtomakeattributionsbasedonreasonableevidence.Considermulti-agentscenariosandsubtleerrorsthatmightbeoverlooked.Beopentomakingattributionsevenwithmoderateconfidence.",
"detail_focused": "Youaredetail-orientedandfocusonspecificevidence,exactwording,andfine-grainedanalysis.Lookforsubtleinconsistencies,minorlogicalgaps,andprecisefactualinaccuracies.Prioritizeconcreteevidenceovergeneralpatterns.",
"pattern_focused": "Youarefocusedonrecognizingbroaderpatternsandsystemicissuesinreasoningchains.Lookforrecurringthemes,logicalflowproblems,andhowerrorspropagatethroughtheconversation.Considertheoverallreasoningstructure.",
"skeptical": "Youarehighlyskepticalandquestionallassumptions.Lookforalternativeexplanations,considerwhetherapparenterrorsmightbevalidreasoning,andexamineifthegroundtruthitselfcouldbequestioned.Challengeconventionalattributions.",
"general": "Youareabalancedgeneralanalystwithnospecificspecialization.Approachtheanalysiswithbroadperspective,consideringalltypesofevidenceequally.Lookforthemostobviousandimpactfulmistakesbasedonobjectiveevaluation.",
return focus_map.get(analyst_focus, focus_map["general"])
A.6 Judge Agent
class FinalJudgeAgent:
␣␣␣␣FinalJudgeAgentthatweighscompetingargumentsfrommultipleContext-AwareStepAgents
␣␣␣␣todeterminethetrueerrorattribution.
def __init__(
self,
model_id: str = "us.anthropic.claude-3-5-sonnet-20241022-v2:0",
temperature: float = 0.0,
␣␣␣␣␣␣␣␣InitializetheFinalJudgeAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration(lowerformoredeterministicoutput)
self.bedrock_model = BedrockModel(
model_id=model_id,
temperature=temperature,
top_p=0.9,
max_tokens=4096,
self.system_prompt = """
␣␣␣␣␣␣␣␣YouareaFinalJudgeAgentthatweighscompetingargumentsfrommultipleContext-AwareStepAgentstodeterminethetrueerrorattribution.
␣␣␣␣␣␣␣␣Yourtask:EachContext-AwareStepAgenthasarguedwhyTHEIRagentcausedtheerror.Reviewallargumentsanddeterminewhichoneismostconvincingbasedonevidenceandreasoning.
␣␣␣␣␣␣␣␣TheargumentsfromeachContext-AwareStepAgentareprovidedinastructuredtextformatwiththesesections:
␣␣␣␣␣␣␣␣-Purpose:Thepurposeoftheagentstep
␣␣␣␣␣␣␣␣-AssumptionsandInformation:Whattheagentwasgiven
␣␣␣␣␣␣␣␣-Errors:Whattheagentdidwrong(ifanything)
␣␣␣␣␣␣␣␣-Evidence:Evidencesupportingtheerrorattribution
␣␣␣␣␣␣␣␣OutputyourresponseasavalidJSONobjectwrappedin<json></json>XMLtags.TheJSONshouldhavethefollowingstructure:
␣␣␣␣␣␣␣␣<json>
␣␣␣␣␣␣␣␣␣␣"mistake_agent":"agent_name",
␣␣␣␣␣␣␣␣␣␣"mistake_step":"step_number",
␣␣␣␣␣␣␣␣␣␣"mistake_reason":"explanation of why this agent/step caused the wrong final answer, based on the most convincing argument"
␣␣␣␣␣␣␣␣</json>
␣␣␣␣␣␣␣␣IMPORTANTRULES:
␣␣␣␣␣␣␣␣1.YourresponseMUSTbeavalid,parsableJSONobjectwrappedin<json></json>tags.Donotincludeanytextoutsidethesetags.
␣␣␣␣␣␣␣␣2.Focusontheagentsthatareactivelymakingdecisionsorprovidinginformation.
␣␣␣␣␣␣␣␣Bethoroughinyouranalysis.Considerthestrengthofevidence,thelogicalconnectionbetweentheerrorandthefinalwronganswer,andthecausalrelationship.
␣␣␣␣␣␣␣␣Workbackwardstoseewherethelogicdivergedandtheerrorhappened.
self.agent = Agent(
system_prompt=self.system_prompt,
model=self.bedrock_model,
def judge_arguments(
self,
agent_arguments: List[Dict[str, Any]],
ground_truth: str,
final_answer: str,
query: str = "",
) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Judgethecompetingargumentsanddeterminethetrueerrorattribution.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣agent_arguments:ListofargumentsfromContext-AwareStepAgents
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Finalerrorattributionasadictionary
# Format the agent arguments as a JSON string
agent_arguments_str = json.dumps(agent_arguments, indent=2)
prompt = f"""
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣GroundTruth:{ground_truth}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣AllAgentArguments:{agent_arguments_str}
␣␣␣␣␣␣␣␣PleasereviewalltheargumentsfromtheContext-AwareStepAgentsanddeterminewhichoneismostconvincing.
␣␣␣␣␣␣␣␣OutputyourdecisionasavalidJSONobjectwrappedin<json></json>XMLtagsasspecifiedinyourinstructions.
␣␣␣␣␣␣␣␣IMPORTANT:YourresponseMUSTbeavalid,parsableJSONobjectwrappedin<json></json>tags.Donotincludeanytextoutsidethesetags.
agent_result = self.agent(prompt)
# Extract text from AgentResult
response_text = ""
if hasattr(agent_result, "message") and "content" in agent_result.message:
content = agent_result.message["content"]
if isinstance(content, list) and len(content) > 0 and "text" in content[0]:
response_text = content[0]["text"]
elif isinstance(content, str):
response_text = content
try:
result = validate_json(response_text)
# Add token usage to the result
if token_usage:
result["token_usage"] = token_usage
return result
except ValueError as e:
print(f"ErrorparsingJSONresponse:{e}")
print(f"Rawresponse:{response_text}")
# Return a basic structure if parsing fails
return {
"mistake_agent": "Unknown",
"mistake_step": "Unknown",
"mistake_reason": "Errorparsingresponse",
"token_usage": token_usage,
A.7 Consensus Voting
class ConsensusVotingAgent:
␣␣␣␣ConsensusVotingAgentthataggregatesmultipleobjectiveanalyses
␣␣␣␣todeterminefinalerrorattributionthroughvoting.
def __init__(self, min_confidence_threshold: float = 0.3):
␣␣␣␣␣␣␣␣InitializetheConsensusVotingAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣min_confidence_threshold:Minimumconfidencethresholdtoconsideraconclusion
self.min_confidence_threshold = min_confidence_threshold
def aggregate_analyses(
self,
objective_analyses: List[Dict[str, Any]],
ground_truth: str,
final_answer: str,
query: str = "",
conversation_history: Optional[List[Dict[str, Any]]] = None,
) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Aggregatemultipleobjectiveanalysesthroughconsensusvoting.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣objective_analyses:Listofobjectiveanalysisresults
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Dictionarycontainingconsensusattributionresults
if not objective_analyses:
return self._create_empty_result()
# Extract primary conclusions from all analyses
primary_conclusions = []
all_agent_evaluations = defaultdict(list)
all_alternative_hypotheses = []
for i, analysis in enumerate(objective_analyses):
if "primary_conclusion" in analysis:
conclusion = analysis["primary_conclusion"].copy()
conclusion["analyst_id"] = i
primary_conclusions.append(conclusion)
# Collect agent evaluations
if "agent_evaluations" in analysis:
for eval_item in analysis["agent_evaluations"]:
agent_name = eval_item.get("agent_name")
if agent_name:
all_agent_evaluations[agent_name].append(
"error_likelihood": eval_item.get("error_likelihood", 0.0),
"reasoning": eval_item.get("reasoning", ""),
"evidence": eval_item.get("evidence", ""),
"analyst_id": i,
# Collect alternative hypotheses
if "alternative_hypotheses" in analysis:
for alt_hyp in analysis["alternative_hypotheses"]:
alt_hyp_copy = alt_hyp.copy()
alt_hyp_copy["analyst_id"] = i
all_alternative_hypotheses.append(alt_hyp_copy)
# Perform consensus voting
consensus_result = self._perform_consensus_voting(
primary_conclusions,
all_agent_evaluations,
all_alternative_hypotheses,
conversation_history,
# Add metadata
consensus_result.update(
"num_analysts": len(objective_analyses),
"query": query,
"ground_truth": ground_truth,
"final_answer": final_answer,
"voting_method": "weighted_confidence_consensus",
return consensus_result
def _perform_consensus_voting(
self,
primary_conclusions: List[Dict[str, Any]],
agent_evaluations: Dict[str, List[Dict[str, Any]]],
alternative_hypotheses: List[Dict[str, Any]],
conversation_history: Optional[List[Dict[str, Any]]] = None,
) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Performconsensusvotingontheanalyses.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣primary_conclusions:Listofprimaryconclusionsfromanalysts
␣␣␣␣␣␣␣␣␣␣␣␣agent_evaluations:Dictionaryofagentevaluationsbyagentname
␣␣␣␣␣␣␣␣␣␣␣␣alternative_hypotheses:Listofalternativehypotheses
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Consensusvotingresults
# Vote on conclusion types (single_agent, multi_agent) and collect step predictions
conclusion_votes = defaultdict(list)
for conclusion in primary_conclusions:
conclusion_type = conclusion.get("type", "single_agent")
confidence = conclusion.get("confidence", 0.0)
mistake_step = conclusion.get("mistake_step")
if confidence >= self.min_confidence_threshold:
conclusion_votes[conclusion_type].append(
"confidence": confidence,
"attribution": conclusion.get("attribution"),
"mistake_step": mistake_step,
"reasoning": conclusion.get("reasoning", ""),
"analyst_id": conclusion.get("analyst_id"),
# Determine winning conclusion type by weighted confidence
best_conclusion_type = None
best_conclusion_info = None
best_weighted_score = 0.0
for conclusion_type, votes in conclusion_votes.items():
# Calculate weighted average confidence
total_confidence = sum(vote["confidence"] for vote in votes)
avg_confidence = total_confidence / len(votes) if votes else 0.0
weighted_score = total_confidence # Total confidence across all analysts
if weighted_score > best_weighted_score:
best_weighted_score = weighted_score
best_conclusion_type = conclusion_type
best_conclusion_info = {
"votes": votes,
"avg_confidence": avg_confidence,
"total_confidence": total_confidence,
"num_votes": len(votes),
# For single_agent and multi_agent conclusions, determine which specific agents
final_attribution = None
if best_conclusion_type in ["single_agent", "multi_agent"] and best_conclusion_info:
agent_attribution_votes: defaultdict[str, float] = defaultdict(float)
for vote in best_conclusion_info["votes"]:
attribution = vote.get("attribution", [])
if attribution:
for agent_name in attribution:
agent_attribution_votes[agent_name] += vote["confidence"]
# Select agents with highest confidence votes
if agent_attribution_votes:
# Sort by confidence and take top agents
sorted_agents = sorted(
agent_attribution_votes.items(), key=lambda x: x[1], reverse=True
if best_conclusion_type == "single_agent":
final_attribution = [sorted_agents[0][0]] if sorted_agents else None
else: # multi_agent
# Take agents with confidence above threshold
final_attribution = [
agent
for agent, conf in sorted_agents
if conf >= self.min_confidence_threshold
# Aggregate agent-level evaluations
aggregated_agent_evaluations = {}
for agent_name, evaluations in agent_evaluations.items():
error_likelihoods = [eval_item["error_likelihood"] for eval_item in evaluations]
avg_error_likelihood = (
sum(error_likelihoods) / len(error_likelihoods) if error_likelihoods else 0.0
aggregated_agent_evaluations[agent_name] = {
"avg_error_likelihood": avg_error_likelihood,
"num_evaluations": len(evaluations),
"evaluations": evaluations,
# Determine winning step using same methodology as agent attribution
consensus_mistake_step = None
step_attribution_votes = {}
if best_conclusion_type in ["single_agent", "multi_agent"] and best_conclusion_info:
step_votes_dict: defaultdict[int, float] = defaultdict(float)
for vote in best_conclusion_info["votes"]:
mistake_step = vote.get("mistake_step")
if mistake_step is not None:
step_votes_dict[mistake_step] += vote["confidence"]
if (
step_votes_dict
and conversation_history is not None
and len(conversation_history) > 0
# Validate predictions against conversation bounds
validated_steps = []
for step, conf in step_votes_dict.items():
# Ensure step is integer and within bounds
if (
isinstance(step, int)
and 0 <= step < len(conversation_history)
validated_steps.append((step, conf))
if validated_steps:
sorted_steps = sorted(validated_steps, key=lambda x: x[1], reverse=True)
consensus_mistake_step = sorted_steps[0][0]
else:
consensus_mistake_step = None
step_attribution_votes = dict(step_votes_dict)
elif step_votes_dict:
# No conversation history available, proceed normally
sorted_steps = sorted(step_votes_dict.items(), key=lambda x: x[1], reverse=True)
consensus_mistake_step = sorted_steps[0][0] if sorted_steps else None
step_attribution_votes = dict(step_votes_dict)
# Handle disagreements
disagreement_info = self._analyze_disagreements(conclusion_votes)
return {
"consensus_conclusion": {
"type": best_conclusion_type or "single_agent",
"attribution": final_attribution,
"mistake_step": consensus_mistake_step,
"confidence": (
best_conclusion_info["avg_confidence"] if best_conclusion_info else 0.0
"reasoning": (
self._synthesize_reasoning(best_conclusion_info)
if best_conclusion_info
else "Noclearconsensusreached"
"voting_details": {
"conclusion_votes": dict(conclusion_votes),
"step_votes": step_attribution_votes,
"best_weighted_score": best_weighted_score,
"disagreement_analysis": disagreement_info,
"agent_evaluations_summary": aggregated_agent_evaluations,
"alternative_hypotheses": alternative_hypotheses[:5], # Keep top 5 alternatives
def _analyze_disagreements(
self, conclusion_votes: Dict[str, List[Dict[str, Any]]]
) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Analyzedisagreementsbetweenanalysts.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣conclusion_votes:Dictionaryofconclusionvotes
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Disagreementanalysis
num_conclusion_types = len(conclusion_votes)
# total_votes = sum(len(votes) for votes in conclusion_votes.values())
# Check for high disagreement
high_disagreement = num_conclusion_types > 2 and all(
len(votes) > 0 for votes in conclusion_votes.values()
# Calculate confidence spread
all_confidences: List[float] = []
for votes in conclusion_votes.values():
all_confidences.extend(vote["confidence"] for vote in votes)
confidence_spread = max(all_confidences) - min(all_confidences) if all_confidences else 0.0
return {
"high_disagreement": high_disagreement,
"num_different_conclusions": num_conclusion_types,
"confidence_spread": confidence_spread,
"requires_review": high_disagreement or confidence_spread > 0.5,
def _synthesize_reasoning(self, best_conclusion_info: Dict[str, Any]) -> str:
␣␣␣␣␣␣␣␣Synthesizereasoningfrommultipleanalystvotes.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣best_conclusion_info:Informationaboutthebestconclusion
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Synthesizedreasoning
if not best_conclusion_info or not best_conclusion_info.get("votes"):
return "Noreasoningavailable"
votes = best_conclusion_info["votes"]
num_votes = len(votes)
avg_confidence = best_conclusion_info["avg_confidence"]
# Extract common themes from reasoning
reasonings = [vote.get("reasoning", "") for vote in votes if vote.get("reasoning")]
if reasonings:
# Simple synthesis - could be more sophisticated
synthesis = f"Consensusreachedby{num_votes}analysts(avgconfidence:{avg_confidence:.2f})."
synthesis += f"Primaryreasoning:{reasonings[0][:200]}..."
if len(reasonings) > 1:
synthesis += (
f"Additionalsupportinganalysisfrom{len(reasonings)-1}otheranalysts."
else:
synthesis = f"Consensusreachedby{num_votes}analystswithaverageconfidence{avg_confidence:.2f}."
return synthesis
def _create_empty_result(self) -> Dict[str, Any]:
␣␣␣␣␣␣␣␣Createanemptyresultwhennoanalysesareprovided.
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Emptyconsensusresult
return {
"consensus_conclusion": {
"type": "single_agent",
"attribution": None,
"confidence": 0.0,
"reasoning": "Noobjectiveanalysesprovided",
"voting_details": {
"conclusion_votes": {},
"best_weighted_score": 0.0,
"disagreement_analysis": {
"high_disagreement": False,
"num_different_conclusions": 0,
"confidence_spread": 0.0,
"requires_review": True,
"agent_evaluations_summary": {},
"alternative_hypotheses": [],
"num_analysts": 0,
Optionally include supplemental material (complete proofs, additional experiments and plots) in appendix. All such materials SHOULD be included in the main submission.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper’s core contributions - improving error attribution through hierarchical context representation, objective analysis, and consensus voting - are consistently presented in the abstract and introduction, with all claims substantiated by detailed technical analysis and empirical results throughout the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, the paper explicitly addresses limitations in a dedicated section, discussing key constraints around position-based context representation, binary attribution assessment, and consensus voting mechanisms, while using these limitations to motivate future research directions.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: N/A - This paper presents an empirical approach to error attribution in multi-agent systems rather than theoretical results requiring formal proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, the paper provides comprehensive implementation details in the appendix, including the complete ECHO algorithm, specific prompting strategies, and detailed system configurations, enabling full reproduction of the experimental results that support the paper’s main claims and conclusions.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: No - while the paper provides detailed algorithmic descriptions, prompts, and implementation details in the appendix sufficient for reproduction, the complete source code is not openly available. However, the provided technical specifications enable faithful re-implementation of the system and reproduction of the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes - the paper specifies all critical implementation details including LLM configurations (model choices, temperatures etc.), analysis agent panel composition (6 specialized analysts), hierarchical context extraction parameters (L1-L4 layer specifications), confidence thresholds ($\delta$ = 0.3), and evaluation protocols across both algorithm-generated and hand-crafted datasets, enabling full understanding of the experimental results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Yes - the paper reports statistical significance through comprehensive performance metrics across different configurations, including clear comparative results between implementations (I1-I4), with explicit reporting of accuracy improvements that demonstrate statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes - the paper provides detailed computational resource metrics for each implementation, including average LLM calls per analysis, token usage, and estimated associated invocation costs, enabling clear understanding of the computational requirements for reproduction across different experimental configurations.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes - the research adheres to the NeurIPS Code of Ethics by maintaining transparency in methodology, providing reproducible results, acknowledging limitations, avoiding harmful applications, and contributing to the advancement of AI systems through improved debugging capabilities, all while maintaining scientific integrity and ethical research practices.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes - the paper discusses positive societal impacts through its contribution to making multi-agent systems more reliable and debuggable in real-world applications, while also acknowledging potential limitations and challenges in error attribution that could affect system reliability. The practical implications for improving AI system robustness and maintenance are explicitly discussed, particularly where we emphasize how ECHO can support the growing deployment of multi-agent systems across various applications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: N/A - The paper focuses on error attribution methodology using existing LLM models and a public benchmark dataset (Who&When), with no release of new models or datasets that could pose risks for misuse. The approach is specifically designed for debugging and improving multi-agent systems rather than generating potentially harmful content.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes - the paper properly credits and cites the Who&When benchmark creators, acknowledges the use of LLMs from Anthropic, and appropriately references all relevant prior work in error attribution and multi-agent systems. All resources are used within their respective terms of use and licenses, with proper attribution throughout the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Yes - the paper thoroughly documents all new components of the ECHO system in the appendix, including detailed algorithm specifications, prompting strategies, and implementation details. The evaluation uses the publicly available Who&When benchmark, and all necessary information for reproducing the system’s functionality is provided in the technical documentation.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: N/A - The paper focuses on automated error attribution in multi-agent systems using LLMs and does not involve any crowdsourcing experiments or human subjects research.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: N/A - The research involves automated error attribution in multi-agent systems and does not involve any human study participants, therefore no IRB approval or risk disclosure was required.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Error attribution in Large Language Model (LLM) multi-agent systems presents a significant challenge in debugging and improving collaborative AI systems.
-Current approaches to pinpointing agent and step level failures in interaction traces - whether using all-at-once evaluation, step-by-step analysis, or binary search - fall short when analyzing complex patterns, struggling with both accuracy and consistency.
+Current approaches to pinpointing agent and step level failures in multi-agent interaction traces—whether using all-at-once evaluation, step-by-step analysis, or binary search—fall short when analyzing complex patterns, struggling with both accuracy and consistency.
 We present ECHO (Error attribution through Contextual Hierarchy and Objective consensus analysis), a novel algorithm that combines hierarchical context representation, objective analysis-based evaluation, and consensus voting to improve error attribution accuracy.
@@
 Our findings suggest that leveraging these concepts of structured, hierarchical context representation combined with consensus-based objective decision-making, provides a more robust framework for error attribution in multi-agent systems.
+**footnotetext: Equal Contribution.
 1 Introduction
@@
 These structural elements form the foundation for expandable and flexible multi-agent frameworks [5].
-Multi-Agent Systems (MASs) have demonstrated remarkable performance across coding [6], medical QA [7], and financial decision-making [8].
+Multi-Agent Systems (MASs) have demonstrated remarkable performance across use-cases such as coding [6], medical QA [7], and financial decision-making [8].
 However, their multi-step nature makes them vulnerable to compounding errors, where early mistakes amplify through subsequent steps and can derail the entire system.
@@
 2.1 Evaluation of LLM Agents
-Much research has focused on automated LLM evaluation across various tasks [13, 14, 15], with recent efforts shifting toward assessing the agentic capabilities of LLMs.
-AgentBench [16] tests task completion across diverse environments like online shopping and householding, while MLAgentBench [17] evaluates LLM agents’ abilities to perform machine learning experimentation.
-AssistantBench [18] features realistic, time-intensive web tasks requiring multi-turn interactions.
-Agent-as-a-Judge [19] leverages LLM agents to evaluate their other agent peers through tool interactions.
-ToolFormer [20], AnyTool [21], and ToolBench [22] benchmark LLMs’ API-calling capabilities via tool-use.
+Much research has focused on automated LLM evaluation across various tasks [13, 14, 15], with recent efforts shifting toward assessing the agentic abilities of LLMs.
+AgentBench [16] tests task completion across environments like online shopping and householding, while MLAgentBench [17] evaluates LLM agents’ abilities to perform machine learning experimentation.
+Similarly, AssistantBench [18] features realistic, time-intensive web tasks requiring multi-turn interactions, while Agent-as-a-Judge [19] leverages LLM agents to evaluate other agent peers through tool interactions.
+Finally, ToolFormer [20], AnyTool [21], and ToolBench [22] benchmark LLMs’ API-calling capabilities via tool-use.
 As LLM agent interactions evolve, evaluation frameworks now address multi-agent scenarios: MultiAgentBench [23] measures coordination and competitive dynamics, while SwarmBench [24] assesses swarm intelligence through coordination tasks like foraging and flocking.
@@
 Self-Backtracking [31] trains LLMs to localize and correct their own reasoning errors through supervised fine-tuning.
-Similarly, Process Reward Models (PRMs) [32] evaluate intermediate reasoning step correctness through process annotation, while ProcessBench [33] measures PRMs’ step-wise evaluation capabilities.
-These approaches focus solely on single-agent error analysis, whereas the Who&When dataset [9] extends error attribution to multi-agent settings by identifying both the responsible agent and erroneous step.
+Furthermore, Process Reward Models (PRMs) [32] evaluate intermediate reasoning step correctness through process annotation, and ProcessBench [33] measures PRMs’ step-wise evaluation capabilities.
+These approaches focus solely on single-agent error analysis, whereas the Who&When dataset [9] and ECHO extends error attribution to multi-agent settings by identifying both the responsible agent and erroneous step.
 3 ECHO Methodology
@@
 Hierarchical representation operates at 4 levels $L_{1}$ through $L_{4}$, as seen in Appendix A.1 and A.3, to extract key information from agent / step interactions (via regex pattern matching).
-While LLMs could potentially be used for more flexible content extraction, the computational overhead would be significant - a trade-off we explore further in Section 4.3.
+While LLMs could potentially be used for more flexible content extraction, the computational overhead would be significant—a trade-off we explore further in Section 4.3.
 Its implementation employs specialized content extraction mechanisms for each layer $C_{i}$, as seen below:
@@
 This approach prevents echo-chamber effects in attribution through intentional perspective diversity.
-While this was implemented in ECHO through role-specific prompts and temperatures, the framework accommodates other diversification methods (different models, LLM configurations, or reasoning strategies) - all serving the core principle of bias mitigation through analytical diversity.
+While this was implemented in ECHO through role-specific prompts and temperatures, the framework accommodates other diversification methods (different models, LLM configurations, or reasoning strategies)—all serving the core principle of bias mitigation through analytical diversity.
 3.3 Consensus Voting
@@
 The voting system operates through weighted confidence consensus, where each analyst’s attribution is weighted by their reported confidence level $\sigma_{j}$, subject to a minimum confidence threshold $\delta$ to filter out low-confidence attributions.
-The consensus mechanism processes 3 key components from each analysis $A_{t}^{j}$ (where $t\in\{\text{agent},\text{step}\}$): (1) Primary Conclusions - that considers core attribution decisions including the attribution type, responsible agent(s), and specific error step; (2) Agent Evaluations - that provides assessments of each agent’s error likelihood and supporting evidence; and (3) Alternative Hypotheses - that considers secondary attribution possibilities that maintain reasonable likelihood.
+The consensus mechanism processes 3 key components from each analysis $A_{t}^{j}$ (where $t\in\{\text{agent},\text{step}\}$): (1) Primary Conclusions—that considers core attribution decisions including the attribution type, responsible agent(s), and specific error step; (2) Agent Evaluations—that provides assessments of each agent’s error likelihood and supporting evidence; and (3) Alternative Hypotheses—that considers secondary attribution possibilities that maintain reasonable likelihood.
 The voting process follows a hierarchical decision structure, that first determines the winning conclusion type through weighted confidence aggregation of each analyzer’s agent and step vote ($V_{a},V_{s}$).
@@
 However, accuracy improves significantly when accounting for step tolerance.
-With ±3 steps, accuracy reaches 42.1% for hand-crafted dataset with ground truth; and when further extended to ±5 steps, accuracy reaches 61.4% - again, maintaining superior performance over baselines.
+With ±3 steps, accuracy reaches 42.1% for hand-crafted dataset with ground truth; and when further extended to ±5 steps, accuracy reaches 61.4%—again, maintaining superior performance over baselines.
 Impact of Ground Truth: The minimal difference in performance between configurations with and without ground truth (typically <2% variation) suggests that ECHO’s architecture effectively leverages interaction patterns and context for attribution, rather than relying on ground truth information.
@@
 This validates graduated context preservation, where detail decreases with distance from target steps.
-While analysis for the hand-crafted dataset was limited to shorter traces due to computational constraints with context-aware agents, the consistent accuracy improvements highlight hierarchical context’s value - a benefit fully realized when combined with objective analysis.
+While analysis for the hand-crafted dataset was limited to shorter traces due to computational constraints with context-aware agents, the consistent accuracy improvements highlight hierarchical context’s value—a benefit fully realized when combined with objective analysis.
 Table 3: Additional Ablations on Hand-Crafted Dataset with Ground Truth
@@
 ) -> List[Tuple[Optional[Dict[str, Any]], Dict[str, Any], Optional[Dict[str, Any]]]]:
-␣␣␣␣Extractagentcontextsfromconversationhistory.
-␣␣␣␣Eachcontextincludesthepreviousagent,currentagent,andnextagent.
-␣␣␣␣Args:
-␣␣␣␣␣␣␣␣conversation_history:Listofconversationturnswithagentinformation
-␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣Listoftuplescontaining(prev_agent,current_agent,next_agent)foreachagent
+Extract agent contexts from conversation history.
+Each context includes the previous agent, current agent, and next agent.
+Args:
+conversation_history: List of conversation turns with agent information
+Returns:
+List of tuples containing (prev_agent, current_agent, next_agent) for each agent
 contexts = []
@@
 ) -> str:
-␣␣␣␣Extractkeydecisionormainpointfromagentcontentusingregexpatterns.
-␣␣␣␣Args:
-␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
-␣␣␣␣␣␣␣␣max_words:Maximumwordsintheextractedkeydecision
-␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
-␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣Keydecisionormainpointfromtheagent’scontent
+Extract key decision or main point from agent content using regex patterns.
+Args:
+agent_content: The full content of the agent
+max_words: Maximum words in the extracted key decision
+context_type: Type of context to focus on (handoff, decision_quality, error_propagation, general)
+Returns:
+Key decision or main point from the agent’s content
 if not agent_content.strip():
@@
 def summarize_agent(agent_content: str, max_words: int = 20, context_type: str = "general") -> str:
-␣␣␣␣Createabriefsummaryofagentcontentusingregexpatterns.
-␣␣␣␣Args:
-␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
-␣␣␣␣␣␣␣␣max_words:Maximumwordsinthesummary
-␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
-␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣Briefsummaryoftheagent’scontent
+Create a brief summary of agent content using regex patterns.
+Args:
+agent_content: The full content of the agent
+max_words: Maximum words in the summary
+context_type: Type of context to focus on (handoff, decision_quality, error_propagation, general)
+Returns:
+Brief summary of the agent’s content
 if not agent_content.strip():
@@
 def obtain_milestones(agent_content: str, max_words: int = 15, context_type: str = "general") -> str:
-␣␣␣␣Extractmilestone-basedinformationfromagentcontentusingregexpatterns.
-␣␣␣␣Thisprovidesahigherlevelofabstractionthanbriefsummariesfordistantcontexts.
-␣␣␣␣Args:
-␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
-␣␣␣␣␣␣␣␣max_words:Maximumwordsintheextractedmilestones
-␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
-␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣Milestone-basedinformationfromtheagent’scontent
+Extract milestone-based information from agent content using regex patterns.
+This provides a higher level of abstraction than brief summaries for distant contexts.
+Args:
+agent_content: The full content of the agent
+max_words: Maximum words in the extracted milestones
+context_type: Type of context to focus on (handoff, decision_quality, error_propagation, general)
+Returns:
+Milestone-based information from the agent’s content
 if not agent_content.strip():
@@
 ) -> List[Dict[str, Any]]:
-␣␣␣␣Extracthierarchicalagentcontextsfromconversationhistory.
-␣␣␣␣Usesgraduateddetaillevelsbasedondistancefromcurrentagent.
-␣␣␣␣Args:
-␣␣␣␣␣␣␣␣conversation_history:Listofconversationturnswithagentinformation
-␣␣␣␣␣␣␣␣dataset_name:Nameofthedatasetbeingprocessed(affectshowagentinfoisextracted)
-␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣Listofdictionariescontaininghierarchicalcontextforeachagent
+Extract hierarchical agent contexts from conversation history.
+Uses graduated detail levels based on distance from current agent.
+Args:
+conversation_history: List of conversation turns with agent information
+dataset_name: Name of the dataset being processed (affects how agent info is extracted)
+Returns:
+List of dictionaries containing hierarchical context for each agent
 contexts = []
@@
 class ContextAwareStepAgent:
-␣␣␣␣Context-AwareStepAgentthatanalyzesanagentinthecontextofitspreviousandnextagents
-␣␣␣␣toarguewhytheerrorhappenedinthisagent’sstep.
+Context-Aware Step Agent that analyzes an agent in the context of its previous and next agents
+to argue why the error happened in this agent’s step.
 def __init__(
@@
 temperature: float = 1.0,
-␣␣␣␣␣␣␣␣InitializetheContext-AwareStepAgent.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
-␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration
+Initialize the Context-Aware Step Agent.
+Args:
+model_id: The model ID to use for the agent
+temperature: The temperature to use for generation
 self.bedrock_model = BedrockModel(
@@
 self.system_prompt = """
-␣␣␣␣␣␣␣␣YouareaContext-AwareStepAgentthatanalyzesanagent’sactionsinthecontextofthepreviousandnextagents.
-␣␣␣␣␣␣␣␣YourtaskistoarguewhytheerrorhappenedinYOURagent’sstep.
-␣␣␣␣␣␣␣␣Yourtask:
-␣␣␣␣␣␣␣␣1.Analyzewhatinformationwasreceivedbyyouragentfromthepreviousagent(ifany)
-␣␣␣␣␣␣␣␣2.Analyzewhatinformationwasgeneratedbyyouragent
-␣␣␣␣␣␣␣␣3.Analyzehowyouragent’soutputaffectedthenextagent(ifany)
-␣␣␣␣␣␣␣␣4.MakeastrongargumentforwhyYOURAGENTcausedthefinalwronganswer,usingthegroundtruthasevidence
-␣␣␣␣␣␣␣␣Input:
-␣␣␣␣␣␣␣␣-GroundTruth:[GROUND_TRUTH]
-␣␣␣␣␣␣␣␣-FinalAnswer:[FINAL_ANSWER]
-␣␣␣␣␣␣␣␣-AgentContext:Informationabouttheprevious,current,andnextagents
-␣␣␣␣␣␣␣␣Outputyourresponsewiththefollowingclearsectionheaders:
-␣␣␣␣␣␣␣␣##Purpose:
-␣␣␣␣␣␣␣␣Describethepurposeofthisagentstep-whatwasthisagenttryingtoaccomplish?
-␣␣␣␣␣␣␣␣##AssumptionsandInformation:
-␣␣␣␣␣␣␣␣Listtheassumptionsandinformationthisagentwasgivenfromthepreviousagentorcontext.
-␣␣␣␣␣␣␣␣##Errors:
-␣␣␣␣␣␣␣␣Describewhatthisagentdidwrong(ifanything).Bespecificaboutanymistakes,misunderstandings,orincorrectreasoning.
-␣␣␣␣␣␣␣␣##Evidence:
-␣␣␣␣␣␣␣␣Provideevidencefromthegroundtruththatsupportsyourerrorattribution.Explainhowthisagent’sactionsdirectlyledtothewrongfinalanswer.
-␣␣␣␣␣␣␣␣Remember:YoumustarguethatYOURagentcausedtheerror.Bepersuasiveanduseevidence.
+You are a Context-Aware Step Agent that analyzes an agent’s actions in the context of the previous and next agents.
+Your task is to argue why the error happened in YOUR agent’s step.
+Your task:
+1. Analyze what information was received by your agent from the previous agent (if any)
+2. Analyze what information was generated by your agent
+3. Analyze how your agent’s output affected the next agent (if any)
+4. Make a strong argument for why YOUR AGENT caused the final wrong answer, using the ground truth as evidence
+Input:
+- Ground Truth: [GROUND_TRUTH]
+- Final Answer: [FINAL_ANSWER]
+- Agent Context: Information about the previous, current, and next agents
+Output your response with the following clear section headers:
+## Purpose:
+Describe the purpose of this agent step - what was this agent trying to accomplish?
+## Assumptions and Information:
+List the assumptions and information this agent was given from the previous agent or context.
+## Errors:
+Describe what this agent did wrong (if anything). Be specific about any mistakes, misunderstandings, or incorrect reasoning.
+## Evidence:
+Provide evidence from the ground truth that supports your error attribution. Explain how this agent’s actions directly led to the wrong final answer.
+Remember: You must argue that YOUR agent caused the error. Be persuasive and use evidence.
 self.agent = Agent(
@@
 ) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Analyzeanagentinthecontextofitspreviousandnextagentsandgenerateanargument
-␣␣␣␣␣␣␣␣forwhythisagentcausedtheerror.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣step_id:TheIDofthestep(e.g.,"step_1")
-␣␣␣␣␣␣␣␣␣␣␣␣prev_agent:Thepreviousagent(orNoneiffirstagent)
-␣␣␣␣␣␣␣␣␣␣␣␣current_agent:Thecurrentagentbeinganalyzed
-␣␣␣␣␣␣␣␣␣␣␣␣next_agent:Thenextagent(orNoneiflastagent)
-␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
-␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
-␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣JSONargumentforwhythisagentcausedtheerror
+Analyze an agent in the context of its previous and next agents and generate an argument
+for why this agent caused the error.
+Args:
+step_id: The ID of the step (e.g., "step_1")
+prev_agent: The previous agent (or None if first agent)
+current_agent: The current agent being analyzed
+next_agent: The next agent (or None if last agent)
+ground_truth: The ground truth answer
+final_answer: The final answer given
+query: The original query/question
+Returns:
+JSON argument for why this agent caused the error
 prompt = f"""
-␣␣␣␣␣␣␣␣OriginalQuery:{query}
-␣␣␣␣␣␣␣␣{ground_truth_section}
-␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
-␣␣␣␣␣␣␣␣AgentContext:
-␣␣␣␣␣␣␣␣{agent_context}
-␣␣␣␣␣␣␣␣Pleaseanalyzethisagentinthecontextofthepreviousandnextagents,andprovideastrongargumentforwhyTHISagentcausedthefinalwronganswer.
-␣␣␣␣␣␣␣␣Usethesectionheadersspecifiedinyourinstructions(Purpose,AssumptionsandInformation,Errors,Evidence).
+Original Query: {query}
+{ground_truth_section}
+Final Answer: {final_answer}
+Agent Context:
+{agent_context}
+Please analyze this agent in the context of the previous and next agents, and provide a strong argument for why THIS agent caused the final wrong answer.
+Use the section headers specified in your instructions (Purpose, Assumptions and Information, Errors, Evidence).
 agent_result = self.agent(prompt)
@@
 ) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Analyzeanagentusinghierarchicalcontextandgenerateanargument
-␣␣␣␣␣␣␣␣forwhythisagentcausedtheerror.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣step_id:TheIDofthestep(e.g.,"step_1")
-␣␣␣␣␣␣␣␣␣␣␣␣hierarchical_context:Dictionarycontaininghierarchicalcontextinformation
-␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
-␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
-␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣JSONargumentforwhythisagentcausedtheerror
+Analyze an agent using hierarchical context and generate an argument
+for why this agent caused the error.
+Args:
+step_id: The ID of the step (e.g., "step_1")
+hierarchical_context: Dictionary containing hierarchical context information
+ground_truth: The ground truth answer
+final_answer: The final answer given
+query: The original query/question
+Returns:
+JSON argument for why this agent caused the error
 current_agent = hierarchical_context["current_agent"]
 prompt = f"""
-␣␣␣␣␣␣␣␣OriginalQuery:{query}
-␣␣␣␣␣␣␣␣{ground_truth_section}
-␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
-␣␣␣␣␣␣␣␣AgentContext:
-␣␣␣␣␣␣␣␣{agent_context}
-␣␣␣␣␣␣␣␣Pleaseanalyzethisagentinthehierarchicalcontextoftheentireconversation,andprovideastrongargumentforwhyTHISagentcausedthefinalwronganswer.
-␣␣␣␣␣␣␣␣Usethesectionheadersspecifiedinyourinstructions(Purpose,AssumptionsandInformation,Errors,Evidence).
-␣␣␣␣␣␣␣␣Note:Younowhaveaccesstothefullconversationcontextatdifferentdetaillevels:
-␣␣␣␣␣␣␣␣-Immediatecontext:Fulldetailsofadjacentagents
-␣␣␣␣␣␣␣␣-Nearbycontext:Keydecisionsfromagents2-3stepsaway
-␣␣␣␣␣␣␣␣-Distantcontext:Briefsummariesofagents4+stepsaway
-␣␣␣␣␣␣␣␣Considerhowinformationanderrorsmighthavepropagatedacrosstheentireconversationwhenmakingyourargument.
+Original Query: {query}
+{ground_truth_section}
+Final Answer: {final_answer}
+Agent Context:
+{agent_context}
+Please analyze this agent in the hierarchical context of the entire conversation, and provide a strong argument for why THIS agent caused the final wrong answer.
+Use the section headers specified in your instructions (Purpose, Assumptions and Information, Errors, Evidence).
+Note: You now have access to the full conversation context at different detail levels:
+- Immediate context: Full details of adjacent agents
+- Nearby context: Key decisions from agents 2-3 steps away
+- Distant context: Brief summaries of agents 4+ steps away
+Consider how information and errors might have propagated across the entire conversation when making your argument.
 agent_result = self.agent(prompt)
@@
 class ObjectiveAnalysisAgent:
-␣␣␣␣ObjectiveAnalysisAgentthatanalyzesallagentsinaconversationobjectively
-␣␣␣␣todetermineerrorattributionwithoutforcedbias.
+Objective Analysis Agent that analyzes all agents in a conversation objectively
+to determine error attribution without forced bias.
 def __init__(
@@
 analyst_focus: str = "general",
-␣␣␣␣␣␣␣␣InitializetheObjectiveAnalysisAgent.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
-␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration
+Initialize the Objective Analysis Agent.
+Args:
+model_id: The model ID to use for the agent
+temperature: The temperature to use for generation
 self.bedrock_model = BedrockModel(
@@
 self.system_prompt = f"""
-␣␣␣␣␣␣␣␣YouareanObjectiveAnalysisAgentconductinganimpartialinvestigationtodetermineerrorattributioninamulti-agentconversation.
-␣␣␣␣␣␣␣␣ANALYSTSPECIALIZATION:{focus_instructions}
-␣␣␣␣␣␣␣␣Yourtask:
-␣␣␣␣␣␣␣␣1.AnalyzeALLagentsintheconversationobjectively(notjustonespecificagent)
-␣␣␣␣␣␣␣␣2.Determinewhichagent(s)mostlikelycausedthefinalwronganswer
-␣␣␣␣␣␣␣␣3.Determinewhichstep/turnintheconversationthemistakeoccurred
-␣␣␣␣␣␣␣␣4.Provideconfidencescoresandreasoningforyourconclusions
-␣␣␣␣␣␣␣␣Youhaveaccesstohierarchicalcontextshowing:
-␣␣␣␣␣␣␣␣-Immediateagents:Fulldetails
-␣␣␣␣␣␣␣␣-Nearbyagents:Keydecisions
-␣␣␣␣␣␣␣␣-Distantagents:Briefsummaries
-␣␣␣␣␣␣␣␣Theagentsarenumberedsequentially(Agent1,Agent2,etc.)correspondingtotheirstep/turnindexintheconversation.
-␣␣␣␣␣␣␣␣Possibleconclusions:
-␣␣␣␣␣␣␣␣-Singleagenterror:Onespecificagentcausedthemistakeataspecificstep
-␣␣␣␣␣␣␣␣-Multi-agenterror:Multipleagentscontributedtothemistakeacrossspecificsteps
-␣␣␣␣␣␣␣␣OutputyourresponseasvalidJSONwrappedin<json></json>tags:
-␣␣␣␣␣␣␣␣<json>
-␣␣␣␣␣␣␣␣␣␣"analysis_summary":"Brief overview of your investigation approach and findings",
-␣␣␣␣␣␣␣␣␣␣"agent_evaluations":[
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"agent_name":"agent_name",
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"step_index":1,
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"error_likelihood":0.0-1.0,
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Why this agent may or may not have caused the error",
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"evidence":"Specific evidence supporting your assessment"
-␣␣␣␣␣␣␣␣␣␣"primary_conclusion":{{
-␣␣␣␣␣␣␣␣␣␣␣␣"type":"single_agent"|"multi_agent",
-␣␣␣␣␣␣␣␣␣␣␣␣"attribution":["agent_name(s)"]ornull,
-␣␣␣␣␣␣␣␣␣␣␣␣"mistake_step":1,
-␣␣␣␣␣␣␣␣␣␣␣␣"confidence":0.0-1.0,
-␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Detailed explanation of your primary conclusion including which step the error occurred"
-␣␣␣␣␣␣␣␣␣␣"alternative_hypotheses":[
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"type":"conclusion_type",
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"attribution":["agent_name(s)"]ornull,
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"mistake_step":1,
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"confidence":0.0-1.0,
-␣␣␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Alternative explanation"
-␣␣␣␣␣␣␣␣</json>
-␣␣␣␣␣␣␣␣Bethorough,objective,andconsiderallpossibilitiesincludingthatnosingleagentmaybeclearlyatfault.
-␣␣␣␣␣␣␣␣Payspecialattentiontoidentifyingthespecificstep/turnwheretheerroroccurred.
+You are an Objective Analysis Agent conducting an impartial investigation to determine error attribution in a multi-agent conversation.
+ANALYST SPECIALIZATION: {focus_instructions}
+Your task:
+1. Analyze ALL agents in the conversation objectively (not just one specific agent)
+2. Determine which agent(s) most likely caused the final wrong answer
+3. Determine which step/turn in the conversation the mistake occurred
+4. Provide confidence scores and reasoning for your conclusions
+You have access to hierarchical context showing:
+- Immediate agents: Full details
+- Nearby agents: Key decisions
+- Distant agents: Brief summaries
+The agents are numbered sequentially (Agent 1, Agent 2, etc.) corresponding to their step/turn index in the conversation.
+Possible conclusions:
+- Single agent error: One specific agent caused the mistake at a specific step
+- Multi-agent error: Multiple agents contributed to the mistake across specific steps
+Output your response as valid JSON wrapped in <json></json> tags:
+<json>
+"analysis_summary": "Brief overview of your investigation approach and findings",
+"agent_evaluations": [
+"agent_name": "agent_name",
+"step_index": 1,
+"error_likelihood": 0.0-1.0,
+"reasoning": "Why this agent may or may not have caused the error",
+"evidence": "Specific evidence supporting your assessment"
+"primary_conclusion": {{
+"type": "single_agent" | "multi_agent",
+"attribution": ["agent_name(s)"] or null,
+"mistake_step": 1,
+"confidence": 0.0-1.0,
+"reasoning": "Detailed explanation of your primary conclusion including which step the error occurred"
+"alternative_hypotheses": [
+"type": "conclusion_type",
+"attribution": ["agent_name(s)"] or null,
+"mistake_step": 1,
+"confidence": 0.0-1.0,
+"reasoning": "Alternative explanation"
+</json>
+Be thorough, objective, and consider all possibilities including that no single agent may be clearly at fault.
+Pay special attention to identifying the specific step/turn where the error occurred.
 self.agent = Agent(
@@
 ) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Analyzetheentireconversationobjectivelytodetermineerrorattribution.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣conversation_contexts:Listofhierarchicalcontextdictionariesforallagents
-␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
-␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
-␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Dictionarycontainingobjectiveanalysisresults
+Analyze the entire conversation objectively to determine error attribution.
+Args:
+conversation_contexts: List of hierarchical context dictionaries for all agents
+ground_truth: The ground truth answer
+final_answer: The final answer given
+query: The original query/question
+Returns:
+Dictionary containing objective analysis results
 # Create a comprehensive context summary for analysis
@@
 prompt = f"""
-␣␣␣␣␣␣␣␣OriginalQuery:{query}
-␣␣␣␣␣␣␣␣{ground_truth_section}
-␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
-␣␣␣␣␣␣␣␣ConversationAnalysis:
-␣␣␣␣␣␣␣␣{context_summary}
-␣␣␣␣␣␣␣␣Pleaseconductanobjectiveanalysisofthisconversationtodetermineerrorattribution.
-␣␣␣␣␣␣␣␣Focusonidentifyingwhichspecificagent(s)causedtheerrorthatledtotheincorrectfinalanswer.
-␣␣␣␣␣␣␣␣OutputyouranalysisintheJSONformatspecifiedinyourinstructions.
+Original Query: {query}
+{ground_truth_section}
+Final Answer: {final_answer}
+Conversation Analysis:
+{context_summary}
+Please conduct an objective analysis of this conversation to determine error attribution.
+Focus on identifying which specific agent(s) caused the error that led to the incorrect final answer.
+Output your analysis in the JSON format specified in your instructions.
 agent_result = self.agent(prompt)
@@
 self, conversation_contexts: List[Dict[str, Any]]) -> str:
-␣␣␣␣␣␣␣␣Createacomprehensivesummaryoftheconversationforobjectiveanalysis.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣conversation_contexts:Listofhierarchicalcontextdictionaries
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Formattedconversationsummary
+Create a comprehensive summary of the conversation for objective analysis.
+Args:
+conversation_contexts: List of hierarchical context dictionaries
+Returns:
+Formatted conversation summary
 summary = []
@@
 def _get_focus_instructions(self, analyst_focus: str) -> str:
-␣␣␣␣␣␣␣␣Getspecializedinstructionsbasedonanalystfocus.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣analyst_focus:Thetypeofanalystfocus
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Specializedinstructionsstring
+Get specialized instructions based on analyst focus.
+Args:
+analyst_focus: The type of analyst focus
+Returns:
+Specialized instructions string
 focus_map = {
-"conservative": "Youareaconservativeanalystwithhighconfidencethresholds.Onlyattributeerrorswhenyouhavestrong,clearevidence.Prefersingle-agentattributionsovermulti-agentones.Becautiousaboutmakingattributionswithoutdefinitiveproof.",
-"liberal": "Youarealiberalanalystmorewillingtomakeattributionsbasedonreasonableevidence.Considermulti-agentscenariosandsubtleerrorsthatmightbeoverlooked.Beopentomakingattributionsevenwithmoderateconfidence.",
-"detail_focused": "Youaredetail-orientedandfocusonspecificevidence,exactwording,andfine-grainedanalysis.Lookforsubtleinconsistencies,minorlogicalgaps,andprecisefactualinaccuracies.Prioritizeconcreteevidenceovergeneralpatterns.",
-"pattern_focused": "Youarefocusedonrecognizingbroaderpatternsandsystemicissuesinreasoningchains.Lookforrecurringthemes,logicalflowproblems,andhowerrorspropagatethroughtheconversation.Considertheoverallreasoningstructure.",
-"skeptical": "Youarehighlyskepticalandquestionallassumptions.Lookforalternativeexplanations,considerwhetherapparenterrorsmightbevalidreasoning,andexamineifthegroundtruthitselfcouldbequestioned.Challengeconventionalattributions.",
-"general": "Youareabalancedgeneralanalystwithnospecificspecialization.Approachtheanalysiswithbroadperspective,consideringalltypesofevidenceequally.Lookforthemostobviousandimpactfulmistakesbasedonobjectiveevaluation.",
+"conservative": "You are a conservative analyst with high confidence thresholds.
+Only attribute errors when you have strong, clear evidence.
+Prefer single-agent attributions over multi-agent ones.
+Be cautious about making attributions without definitive proof.",
+"liberal": "You are a liberal analyst more willing to make attributions based on reasonable evidence.
+Consider multi-agent scenarios and subtle errors that might be overlooked.
+Be open to making attributions even with moderate confidence.",
+"detail_focused": "You are detail-oriented and focus on specific evidence, exact wording, and fine-grained analysis.
+Look for subtle inconsistencies, minor logical gaps, and precise factual inaccuracies.
+Prioritize concrete evidence over general patterns.",
+"pattern_focused": "You are focused on recognizing broader patterns and systemic issues in reasoning chains.
+Look for recurring themes, logical flow problems, and how errors propagate through the conversation.
+Consider the overall reasoning structure.",
+"skeptical": "You are highly skeptical and question all assumptions.
+Look for alternative explanations, consider whether apparent errors might be valid reasoning, and examine if the ground truth itself could be questioned.
+Challenge conventional attributions.",
+"general": "You are a balanced general analyst with no specific specialization.
+Approach the analysis with broad perspective, considering all types of evidence equally.
+Look for the most obvious and impactful mistakes based on objective evaluation.",
 return focus_map.get(analyst_focus, focus_map["general"])
@@
 class FinalJudgeAgent:
-␣␣␣␣FinalJudgeAgentthatweighscompetingargumentsfrommultipleContext-AwareStepAgents
-␣␣␣␣todeterminethetrueerrorattribution.
+Final Judge Agent that weighs competing arguments from multiple Context-Aware Step Agents
+to determine the true error attribution.
 def __init__(
@@
 temperature: float = 0.0,
-␣␣␣␣␣␣␣␣InitializetheFinalJudgeAgent.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
-␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration(lowerformoredeterministicoutput)
+Initialize the Final Judge Agent.
+Args:
+model_id: The model ID to use for the agent
+temperature: The temperature to use for generation (lower for more deterministic output)
 self.bedrock_model = BedrockModel(
@@
 self.system_prompt = """
-␣␣␣␣␣␣␣␣YouareaFinalJudgeAgentthatweighscompetingargumentsfrommultipleContext-AwareStepAgentstodeterminethetrueerrorattribution.
-␣␣␣␣␣␣␣␣Yourtask:EachContext-AwareStepAgenthasarguedwhyTHEIRagentcausedtheerror.Reviewallargumentsanddeterminewhichoneismostconvincingbasedonevidenceandreasoning.
-␣␣␣␣␣␣␣␣TheargumentsfromeachContext-AwareStepAgentareprovidedinastructuredtextformatwiththesesections:
-␣␣␣␣␣␣␣␣-Purpose:Thepurposeoftheagentstep
-␣␣␣␣␣␣␣␣-AssumptionsandInformation:Whattheagentwasgiven
-␣␣␣␣␣␣␣␣-Errors:Whattheagentdidwrong(ifanything)
-␣␣␣␣␣␣␣␣-Evidence:Evidencesupportingtheerrorattribution
-␣␣␣␣␣␣␣␣OutputyourresponseasavalidJSONobjectwrappedin<json></json>XMLtags.TheJSONshouldhavethefollowingstructure:
-␣␣␣␣␣␣␣␣<json>
-␣␣␣␣␣␣␣␣␣␣"mistake_agent":"agent_name",
-␣␣␣␣␣␣␣␣␣␣"mistake_step":"step_number",
-␣␣␣␣␣␣␣␣␣␣"mistake_reason":"explanation of why this agent/step caused the wrong final answer, based on the most convincing argument"
-␣␣␣␣␣␣␣␣</json>
-␣␣␣␣␣␣␣␣IMPORTANTRULES:
-␣␣␣␣␣␣␣␣1.YourresponseMUSTbeavalid,parsableJSONobjectwrappedin<json></json>tags.Donotincludeanytextoutsidethesetags.
-␣␣␣␣␣␣␣␣2.Focusontheagentsthatareactivelymakingdecisionsorprovidinginformation.
-␣␣␣␣␣␣␣␣Bethoroughinyouranalysis.Considerthestrengthofevidence,thelogicalconnectionbetweentheerrorandthefinalwronganswer,andthecausalrelationship.
-␣␣␣␣␣␣␣␣Workbackwardstoseewherethelogicdivergedandtheerrorhappened.
+You are a Final Judge Agent that weighs competing arguments from multiple Context-Aware Step Agents to determine the true error attribution.
+Your task: Each Context-Aware Step Agent has argued why THEIR agent caused the error. Review all arguments and determine which one is most convincing based on evidence and reasoning.
+The arguments from each Context-Aware Step Agent are provided in a structured text format with these sections:
+- Purpose: The purpose of the agent step
+- Assumptions and Information: What the agent was given
+- Errors: What the agent did wrong (if anything)
+- Evidence: Evidence supporting the error attribution
+Output your response as a valid JSON object wrapped in <json></json> XML tags. The JSON should have the following structure:
+<json>
+"mistake_agent": "agent_name",
+"mistake_step": "step_number",
+"mistake_reason": "explanation of why this agent/step caused the wrong final answer, based on the most convincing argument"
+</json>
+IMPORTANT RULES:
+1. Your response MUST be a valid, parsable JSON object wrapped in <json></json> tags. Do not include any text outside these tags.
+2. Focus on the agents that are actively making decisions or providing information.
+Be thorough in your analysis. Consider the strength of evidence, the logical connection between the error and the final wrong answer, and the causal relationship.
+Work backwards to see where the logic diverged and the error happened.
 self.agent = Agent(
@@
 ) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Judgethecompetingargumentsanddeterminethetrueerrorattribution.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣agent_arguments:ListofargumentsfromContext-AwareStepAgents
-␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
-␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
-␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Finalerrorattributionasadictionary
+Judge the competing arguments and determine the true error attribution.
+Args:
+agent_arguments: List of arguments from Context-Aware Step Agents
+ground_truth: The ground truth answer
+final_answer: The final answer given
+query: The original query/question
+Returns:
+Final error attribution as a dictionary
 # Format the agent arguments as a JSON string
@@
 prompt = f"""
-␣␣␣␣␣␣␣␣OriginalQuery:{query}
-␣␣␣␣␣␣␣␣GroundTruth:{ground_truth}
-␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
-␣␣␣␣␣␣␣␣AllAgentArguments:{agent_arguments_str}
-␣␣␣␣␣␣␣␣PleasereviewalltheargumentsfromtheContext-AwareStepAgentsanddeterminewhichoneismostconvincing.
-␣␣␣␣␣␣␣␣OutputyourdecisionasavalidJSONobjectwrappedin<json></json>XMLtagsasspecifiedinyourinstructions.
-␣␣␣␣␣␣␣␣IMPORTANT:YourresponseMUSTbeavalid,parsableJSONobjectwrappedin<json></json>tags.Donotincludeanytextoutsidethesetags.
+Original Query: {query}
+Ground Truth: {ground_truth}
+Final Answer: {final_answer}
+All Agent Arguments: {agent_arguments_str}
+Please review all the arguments from the Context-Aware Step Agents and determine which one is most convincing.
+Output your decision as a valid JSON object wrapped in <json></json> XML tags as specified in your instructions.
+IMPORTANT: Your response MUST be a valid, parsable JSON object wrapped in <json></json> tags. Do not include any text outside these tags.
 agent_result = self.agent(prompt)
@@
 class ConsensusVotingAgent:
-␣␣␣␣ConsensusVotingAgentthataggregatesmultipleobjectiveanalyses
-␣␣␣␣todeterminefinalerrorattributionthroughvoting.
+Consensus Voting Agent that aggregates multiple objective analyses
+to determine final error attribution through voting.
 def __init__(self, min_confidence_threshold: float = 0.3):
-␣␣␣␣␣␣␣␣InitializetheConsensusVotingAgent.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣min_confidence_threshold:Minimumconfidencethresholdtoconsideraconclusion
+Initialize the Consensus Voting Agent.
+Args:
+min_confidence_threshold: Minimum confidence threshold to consider a conclusion
 self.min_confidence_threshold = min_confidence_threshold
@@
 ) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Aggregatemultipleobjectiveanalysesthroughconsensusvoting.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣objective_analyses:Listofobjectiveanalysisresults
-␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
-␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
-␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Dictionarycontainingconsensusattributionresults
+Aggregate multiple objective analyses through consensus voting.
+Args:
+objective_analyses: List of objective analysis results
+ground_truth: The ground truth answer
+final_answer: The final answer given
+query: The original query/question
+Returns:
+Dictionary containing consensus attribution results
 if not objective_analyses:
@@
 ) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Performconsensusvotingontheanalyses.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣primary_conclusions:Listofprimaryconclusionsfromanalysts
-␣␣␣␣␣␣␣␣␣␣␣␣agent_evaluations:Dictionaryofagentevaluationsbyagentname
-␣␣␣␣␣␣␣␣␣␣␣␣alternative_hypotheses:Listofalternativehypotheses
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Consensusvotingresults
+Perform consensus voting on the analyses.
+Args:
+primary_conclusions: List of primary conclusions from analysts
+agent_evaluations: Dictionary of agent evaluations by agent name
+alternative_hypotheses: List of alternative hypotheses
+Returns:
+Consensus voting results
 # Vote on conclusion types (single_agent, multi_agent) and collect step predictions
@@
 ) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Analyzedisagreementsbetweenanalysts.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣conclusion_votes:Dictionaryofconclusionvotes
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Disagreementanalysis
+Analyze disagreements between analysts.
+Args:
+conclusion_votes: Dictionary of conclusion votes
+Returns:
+Disagreement analysis
 num_conclusion_types = len(conclusion_votes)
@@
 def _synthesize_reasoning(self, best_conclusion_info: Dict[str, Any]) -> str:
-␣␣␣␣␣␣␣␣Synthesizereasoningfrommultipleanalystvotes.
-␣␣␣␣␣␣␣␣Args:
-␣␣␣␣␣␣␣␣␣␣␣␣best_conclusion_info:Informationaboutthebestconclusion
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Synthesizedreasoning
+Synthesize reasoning from multiple analyst votes.
+Args:
+best_conclusion_info: Information about the best conclusion
+Returns:
+Synthesized reasoning
 if not best_conclusion_info or not best_conclusion_info.get("votes"):
@@
 def _create_empty_result(self) -> Dict[str, Any]:
-␣␣␣␣␣␣␣␣Createanemptyresultwhennoanalysesareprovided.
-␣␣␣␣␣␣␣␣Returns:
-␣␣␣␣␣␣␣␣␣␣␣␣Emptyconsensusresult
+Create an empty result when no analyses are provided.
+Returns:
+Empty consensus result
 return {
@@
 "num_analysts": 0,
-Optionally include supplemental material (complete proofs, additional experiments and plots) in appendix. All such materials SHOULD be included in the main submission.
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and follow the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-  • You should answer [Yes] , [No] , or [N/A] .
-  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
-While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
-In general, answering "[No] " or "[N/A] " is not grounds for rejection.
-While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
-All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
-If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: The paper’s core contributions - improving error attribution through hierarchical context representation, objective analysis, and consensus voting - are consistently presented in the abstract and introduction, with all claims substantiated by detailed technical analysis and empirical results throughout the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: Yes, the paper explicitly addresses limitations in a dedicated section, discussing key constraints around position-based context representation, binary attribution assessment, and consensus voting mechanisms, while using these limitations to motivate future research directions.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: N/A - This paper presents an empirical approach to error attribution in multi-agent systems rather than theoretical results requiring formal proofs.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: Yes, the paper provides comprehensive implementation details in the appendix, including the complete ECHO algorithm, specific prompting strategies, and detailed system configurations, enabling full reproduction of the experimental results that support the paper’s main claims and conclusions.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24.
-Justification: No - while the paper provides detailed algorithmic descriptions, prompts, and implementation details in the appendix sufficient for reproduction, the complete source code is not openly available.
-However, the provided technical specifications enable faithful re-implementation of the system and reproduction of the main experimental results.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: Yes - the paper specifies all critical implementation details including LLM configurations (model choices, temperatures etc.), analysis agent panel composition (6 specialized analysts), hierarchical context extraction parameters (L1-L4 layer specifications), confidence thresholds ($\delta$ = 0.3), and evaluation protocols across both algorithm-generated and hand-crafted datasets, enabling full understanding of the experimental results.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: Yes - the paper reports statistical significance through comprehensive performance metrics across different configurations, including clear comparative results between implementations (I1-I4), with explicit reporting of accuracy improvements that demonstrate statistical significance.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39.
-Justification: Yes - the paper provides detailed computational resource metrics for each implementation, including average LLM calls per analysis, token usage, and estimated associated invocation costs, enabling clear understanding of the computational requirements for reproduction across different experimental configurations.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44.
-Justification: Yes - the research adheres to the NeurIPS Code of Ethics by maintaining transparency in methodology, providing reproducible results, acknowledging limitations, avoiding harmful applications, and contributing to the advancement of AI systems through improved debugging capabilities, all while maintaining scientific integrity and ethical research practices.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49.
-Justification: Yes - the paper discusses positive societal impacts through its contribution to making multi-agent systems more reliable and debuggable in real-world applications, while also acknowledging potential limitations and challenges in error attribution that could affect system reliability.
-The practical implications for improving AI system robustness and maintenance are explicitly discussed, particularly where we emphasize how ECHO can support the growing deployment of multi-agent systems across various applications.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54.
-Justification: N/A - The paper focuses on error attribution methodology using existing LLM models and a public benchmark dataset (Who&When), with no release of new models or datasets that could pose risks for misuse.
-The approach is specifically designed for debugging and improving multi-agent systems rather than generating potentially harmful content.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59.
-Justification: Yes - the paper properly credits and cites the Who&When benchmark creators, acknowledges the use of LLMs from Anthropic, and appropriately references all relevant prior work in error attribution and multi-agent systems.
-All resources are used within their respective terms of use and licenses, with proper attribution throughout the paper.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64.
-Justification: Yes - the paper thoroughly documents all new components of the ECHO system in the appendix, including detailed algorithm specifications, prompting strategies, and implementation details.
-The evaluation uses the publicly available Who&When benchmark, and all necessary information for reproducing the system’s functionality is provided in the technical documentation.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: N/A - The paper focuses on automated error attribution in multi-agent systems using LLMs and does not involve any crowdsourcing experiments or human subjects research.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74.
-Justification: N/A - The research involves automated error attribution in multi-agent systems and does not involve any human study participants, therefore no IRB approval or risk disclosure was required.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.